TREX: Logbook

TREX: Chronological Logbook Entries

Return to Logbook Top Page
Entry Date Title Site Author #Graphics
268 Mon 01-May-2006West data copiedwestpoulos
267 Mon 01-May-2006South data copiedsouthpoulos
266 Mon 01-May-2006Final status and tear down duties in briefallpoulos
265 Sun 30-Apr-2006Inverter failure ends hot-film run at 1000 LT 30 Aprcentralpoulos
264 Sun 30-Apr-2006Lent gravimetric/TRIME to Berkeley/Stanford de-install teamallpoulos
263 Sun 30-Apr-2006Daily statusallpoulos
262 Sun 30-Apr-2006H2 levelcentralgolubieski
261 Sat 29-Apr-2006Trime Measurements, etc.allgolubieski
260 Sat 29-Apr-2006Guy Wire Tensionswestgolubieski
259 Sat 29-Apr-2006Guy Wire Tensionssouthgolubieski
258 Sat 29-Apr-2006ISFF media daycentralpoulos
257 Sat 29-Apr-2006Daily statusallpoulos
256 Fri 28-Apr-2006Central Guy Tensionscentralgolubieski
255 Fri 28-Apr-2006Krypton Cleaningcentralgolubieski
254 Fri 28-Apr-2006Daily statusallpoulos
253 Thu 27-Apr-2006Krypton cleanings - South, Central, Westallpoulos
252 Thu 27-Apr-2006T/RH 30m West restartedwestpoulos
251 Thu 27-Apr-2006ISFF T-REX daily status 27 Apr 2006allpoulos
250 Wed 26-Apr-2006Redundant copies for exthd7, 8, 9, 10allpoulos
249 Wed 26-Apr-2006Radiation Sensor Cleaningallagee
248 Wed 26-Apr-2006Soil Sensor Photosallagee
247 Wed 26-Apr-2006Trime Measurements, Gravimetricsouthagee
246 Wed 26-Apr-2006eantE DLR lidar Etherant returnedbasepoulos
245 Wed 26-Apr-2006Aster outage when docking station hardware suddenly and permanently failedbasepoulos
244 Wed 26-Apr-2006Daily statusallpoulos
243 Tue 25-Apr-2006Wireless fiddle testscentralpoulos
242 Tue 25-Apr-2006Wireless intermittent outages, return to the original configuration 12.40 centralpoulos
241 Tue 25-Apr-2006Daily status 26 Apr 2006allpoulos
240 Mon 24-Apr-2006Hotfilm runcentralmilitzer
239 Mon 24-Apr-2006Cleaned 5m Kryptoncentralmilitzer
238 Mon 24-Apr-2006Rain Check, Download Datacentralmilitzer
237 Mon 24-Apr-2006Tsoil Coefficients, etc.centralmilitzer
236 Mon 24-Apr-2006Trime Measurements, Gravimetriccentralmilitzer
235 Mon 24-Apr-2006Trime Measurement, Gravimetricwestagee
234 Mon 24-Apr-20065m, 30m Kryptons cleanedwestmilitzer
233 Mon 24-Apr-2006Daily statusallmilitzer
232 Sun 23-Apr-2006Hydrogen replaced at Centralcentralpoulos
231 Sun 23-Apr-2006Rebooted TPOP, Slight adjust of EantJ on centralallmilitzer
230 Sun 23-Apr-2006changed network route to westbasemaclean
229 Sun 23-Apr-2006Daily statusallpoulos
228 Sat 22-Apr-2006Boom angles, centralcentralpoulos
227 Sat 22-Apr-2006Guy wire tensionssouthpoulos
226 Sat 22-Apr-20065m, 30m Kryptons cleanedsouthmilitzer
225 Sat 22-Apr-2006Outages continue; New Eant at Centralcentralmilitzer
224 Fri 21-Apr-2006Aerosol Laser Bad, return to factorywestmilitzer
223 Fri 21-Apr-2006GPS coordinates of the hot-film sonic anemometerscentralpoulos
222 Fri 21-Apr-2006Pictures around Licorcentralpoulos
221 Fri 21-Apr-2006Central pocketec downloadcentralpoulos
220 Fri 21-Apr-2006Daily statusallpoulos
219 Thu 20-Apr-2006hotfilm runningcentralmilitzer
218 Thu 20-Apr-20065m Krypton cleanedcentralmilitzer
217 Thu 20-Apr-2006Trime Measurement, Gravimetricsouthmilitzer
216 Thu 20-Apr-2006Rad.cleanedsouthmilitzer
215 Thu 20-Apr-2006Daily statusallmilitzer
214 Wed 19-Apr-2006Aerosol downwestmilitzer
213 Wed 19-Apr-2006Trime Measurement, Gravimetriccentralmilitzer
212 Wed 19-Apr-2006Rad.cleanedcentralmilitzer
211 Wed 19-Apr-2006Daily statusallpoulos
210 Tue 18-Apr-2006Trime Measurement, Gravimetricwestmilitzer
209 Tue 18-Apr-2006Rad.cleaned, Rnet leveledwestmilitzer
208 Tue 18-Apr-2006Installed Charge-Monitor, Rebootwestmilitzer
207 Tue 18-Apr-2006Boom angles shotwestpoulos
206 Tue 18-Apr-2006TP01 wiring, possible fixwestmilitzer
205 Tue 18-Apr-2006West guy wire tension check & re-tensionwestpoulos
204 Tue 18-Apr-2006hotfilm system clock aheadcentralmaclean
203 Tue 18-Apr-2006Daily statusallpoulos
202 Mon 17-Apr-2006Daily statusallmilitzer
201 Sun 16-Apr-2006Started hotfilmcentralmaclean
200 Sun 16-Apr-2006Daily statusallpoulos
198 Sat 15-Apr-2006Boom anglessouthpoulos
197 Sat 15-Apr-2006hotfilm powercentralmaclean
196 Sat 15-Apr-2006Short local storage outagecentralmilitzer
195 Sat 15-Apr-2006Reconfigured power to hotfilm tower.centralmilitzer
194 Sat 15-Apr-200610m Rad cleaned, photoscentralpoulos
193 Sat 15-Apr-20065m TRH filter checkcentralmilitzer
192 Sat 15-Apr-2006Daily statusallmilitzer
191 Fri 14-Apr-2006hotfilm data system testcentralpoulos
190 Fri 14-Apr-2006Central 2m sonic, power usagecentralpoulos
189 Fri 14-Apr-2006Hacker Attack - Changed ssh daemon allowsmilitzer
188 Fri 14-Apr-2006Local storage retrievedwestmilitzer
187 Fri 14-Apr-2006Daily statusallmilitzer
186 Thu 13-Apr-2006hotfilm data system testcentralmaclean
185 Thu 13-Apr-20062m Sonic downcentralmaclean
184 Thu 13-Apr-2006Updates on Echo/Qsoil negative at southsouthhorst
182 Thu 13-Apr-2006Update on 35W Central Fuel Cell centralmilitzer
181 Thu 13-Apr-2006Guy Wires added to 5m towercentralmilitzer
180 Thu 13-Apr-200630m krypton cleanedcentralmilitzer
179 Thu 13-Apr-2006Daily statusallmilitzer
178 Wed 12-Apr-2006Update on Central Fuel Cell / Battery Statuscentralmilitzer
177 Wed 12-Apr-2006'cleaned' 10m K&Zcentralmilitzer
176 Wed 12-Apr-20065m krypton cleanedcentralmilitzer
175 Wed 12-Apr-2006Daily statusallmilitzer
174 Tue 11-Apr-2006Update on Central Fuel Cell / Battery Statuscentralmilitzer
173 Tue 11-Apr-2006Removed network equipment from schoolbasemaclean
172 Tue 11-Apr-2006Daily statusallmilitzer
171 Mon 10-Apr-2006Hotfilm runcentralmaclean
170 Mon 10-Apr-2006restarted data systemsallmaclean
169 Mon 10-Apr-2006Daily statusallmilitzer
168 Sun 09-Apr-2006Update on Central Fuel Cell / Battery Statuscentralmilitzer
167 Sun 09-Apr-2006guy wire tensionsallmilitzer
166 Sun 09-Apr-2006Radiometers 'Cleaned'westmilitzer
165 Sun 09-Apr-2006Radiometers 'Cleaned'centralmilitzer
164 Sun 09-Apr-2006Radiometers 'Cleaned'southmilitzer
163 Sun 09-Apr-2006gravimetric soil moisturesouthmilitzer
162 Sun 09-Apr-2006Daily statusallmilitzer
161 Sun 09-Apr-2006HotFilm Run Sa. 4/8/06 - Su. 4/9/06centralmilitzer
160 Fri 07-Apr-2006gravimetric soil moisturecentralmilitzer
159 Fri 07-Apr-2006Central Battery Statuscentralmilitzer
158 Fri 07-Apr-2006Imitation Hotfilm Test Runningcentralmilitzer
157 Fri 07-Apr-2006Hotfilm Shutdowncentralmilitzer
156 Thu 06-Apr-2006gravimetric soil moisturewesthorst
155 Thu 06-Apr-2006no rain data at centralcentralhorst
154 Thu 06-Apr-2006wireless network from central downcentralhorst
153 Thu 06-Apr-2006Daily statusallhorst
152 Wed 05-Apr-2006Hotfilm runningcentralsemmer
151 Wed 05-Apr-2006central communications lostcentralgolubieski
150 Tue 04-Apr-2006ZigBee now power on/off at Centralcentralsemmer
149 Tue 04-Apr-2006missing data from logger at westwesthorst
148 Tue 04-Apr-2006Daily statusallhorst
147 Tue 04-Apr-2006lost rf connection to centralcentralhorst
146 Tue 04-Apr-2006loggers at West and Southsouthsemmer
145 Tue 04-Apr-2006Link to central downcentralsemmer
144 Mon 03-Apr-2006Intermittent logger datasouthhorst
143 Mon 03-Apr-2006replaced MaxStream radio in logger at southsouthhorst
142 Mon 03-Apr-2006retensioned guy wirescentralhorst
141 Mon 03-Apr-2006hot film procedurescentralsemmer
140 Mon 03-Apr-2006Daily statusallhorst
139 Sun 02-Apr-2006Daily statusallhorst
138 Sun 02-Apr-2006how to restart covarsbasehorst
137 Sun 02-Apr-2006Krypton clean and s/n at Westwestsemmer
136 Sun 02-Apr-2006Hot film statuscentralsemmer
135 Sun 02-Apr-2006Rad system ar west runningsemmer
134 Sat 01-Apr-200625 m sonic at centralcentralhorst
133 Sat 01-Apr-2006restared wireless network and dsm serverbasehorst
132 Sat 01-Apr-2006maintenance at centralcentralhorst
131 Fri 31-Mar-2006logger at westwestsemmer
130 Fri 31-Mar-2006Daily statusallhorst
129 Thu 30-Mar-2006Krypton cals & locationallsemmer
128 Thu 30-Mar-2006gravimetric soil moistureallhorst
127 Thu 30-Mar-2006HotFilm shutoffcentralsemmer
126 Thu 30-Mar-2006Qsoil negative at southsouthhorst
125 Thu 30-Mar-2006Daily statusallhorst
124 Wed 29-Mar-2006Hot Film statuscentralsemmer
123 Wed 29-Mar-2006Logger #1 at centralcentralsemmer
122 Wed 29-Mar-2006guy wire tensionsallhorst
121 Wed 29-Mar-2006Shutdown pyg sensor at central, logger ID 2centralsemmer
120 Wed 29-Mar-2006Daily statusallhorst
119 Tue 28-Mar-2006Hot film statuscentralsemmer
118 Tue 28-Mar-2006Hot film sonics on central nidscentralsemmer
117 Tue 28-Mar-2006hot-film sonics added to NIDScentralhorst
116 Tue 28-Mar-2006Daily statusallhorst
115 Mon 27-Mar-2006HF site back up collecting sonic data at 15:30centralsemmer
114 Mon 27-Mar-2006Copy local storage to aster over the networkwesthorst
113 Mon 27-Mar-2006Battery changed at logger #1 at centralcentralsemmer
112 Mon 27-Mar-2006Hot film statuscentralsemmer
111 Mon 27-Mar-2006Daily statusallhorst
110 Sun 26-Mar-2006Deploying hot film probescentralhorst
109 Sun 26-Mar-2006Daily statusallhorst
108 Sun 26-Mar-2006dsm server downbasehorst
107 Sun 26-Mar-2006PAM mast broke last nightbasehorst
106 Sat 25-Mar-2006hotfilm updatecentralsemmer
105 Sat 25-Mar-2006replaced fence battery at west towerwesthorst
104 Sat 25-Mar-2006Daily statusallhorst
103 Fri 24-Mar-2006hotfilm statuscentralsemmer
102 Fri 24-Mar-2006Cleaned KH2O at westwesthorst
101 Fri 24-Mar-2006data link to central fixedcentralhorst
100 Fri 24-Mar-2006Daily statusallhorst
99 Fri 24-Mar-2006Hotfilm work from yesterdaycentralsemmer
98 Thu 23-Mar-2006maintenance at west sitewesthorst
97 Thu 23-Mar-2006maintenance at central sitecentralhorst
96 Thu 23-Mar-2006Maintenance at south yesterdaysouthhorst
95 Thu 23-Mar-2006Daily statusallhorst
94 Thu 23-Mar-2006hotfilmcentralsemmer
93 Wed 22-Mar-2006Network disconnectedbasesemmer
92 Wed 22-Mar-2006CRASH!basesemmer
91 Wed 22-Mar-2006htofilm - probe position problemcentralsemmer
90 Wed 22-Mar-2006Installed hot film sensorscentralhorst
89 Wed 22-Mar-2006Cleaning KH2O windowsallhorst
88 Wed 22-Mar-2006Daily statusallhorst
87 Tue 21-Mar-2006hotfilmcentralsemmer
86 Tue 21-Mar-2006hotfilm statuscentralsemmer
85 Tue 21-Mar-2006Daily statusallhorst
84 Mon 20-Mar-2006How to restart the dsm_serverbasesemmer
83 Mon 20-Mar-2006Htofilm statuscentralsemmer
82 Mon 20-Mar-2006Recycled power at schoolhorst
81 Mon 20-Mar-2006KH2O serializer dataallhorst
80 Mon 20-Mar-2006Daily statusallhorst
79 Mon 20-Mar-2006hotfilmcentralsemmer
78 Sun 19-Mar-2006fuel cellcentralsemmer
77 Sun 19-Mar-2006hotfilm statuscentralsemmer
76 Sun 19-Mar-2006Daily statusallhorst
75 Sun 19-Mar-2006HotFilm statusbasesemmer
74 Sat 18-Mar-2006central fuel cellcentralsemmer
73 Sat 18-Mar-2006Central visitcentralsemmer
72 Sat 18-Mar-2006Daily statusallhorst
71 Sat 18-Mar-2006Fuel Cell at centralcentralsemmer
70 Sat 18-Mar-2006hot film run in basebasesemmer
69 Fri 17-Mar-2006HotFilmcentralsemmer
68 Fri 17-Mar-2006Removed hot film data system and 1 m sonic centralhorst
67 Fri 17-Mar-200660 watt fuel cellcentralsemmer
66 Fri 17-Mar-2006single pyg unitscentralsemmer
65 Fri 17-Mar-2006Daily statusallhorst
64 Thu 16-Mar-2006west visitwestoncley
63 Thu 16-Mar-2006South visitsouthoncley
62 Thu 16-Mar-2006Central visitcentraloncley
61 Thu 16-Mar-2006one-time tasksalloncley
60 Thu 16-Mar-2006gravimetric soil measurementsalloncley
59 Thu 16-Mar-2006prometheus config changecentraloncley
58 Thu 16-Mar-2006Daily statusallhorst
57 Wed 15-Mar-2006Service visit to Central towercentralhorst
56 Wed 15-Mar-2006Service visit to West towerwesthorst
55 Wed 15-Mar-2006Service visit to South towersouthhorst
54 Wed 15-Mar-2006hotfilm analogs diedcentraloncley
53 Tue 14-Mar-2006cylinder readingscentraloncley
52 Tue 14-Mar-200615 m TRH failed at south sitesouthhorst
51 Tue 14-Mar-2006TP01 wiring fixedwestoncley
50 Tue 14-Mar-2006Tcase working at centralcentraloncley
49 Tue 14-Mar-2006South service visitsouthhorst
48 Tue 14-Mar-2006West service visitwesthorst
47 Mon wiring errorcentraloncley
46 Mon 13-Mar-2006Daily statusalloncley
45 Sun 12-Mar-2006central died last nightcentraloncley
44 Sun 12-Mar-2006south visitsouthagee
43 Sun 12-Mar-2006hotfilm stuff installedcentraloncley
42 Sun 12-Mar-2006west visitwestoncley
41 Sat 11-Mar-2006Daily statusalloncley
40 Sat 11-Mar-2006visitwestoncley
39 Sat 11-Mar-2006visitcentraloncley
38 Sat 11-Mar-2006added fuel cellcentraloncley
37 Fri 10-Mar-2006Daily statusalloncley
36 Fri 10-Mar-2006Radiometer Power Issuescentralagee
35 Fri 10-Mar-2006dsm crashbaseoncley
34 Thu 09-Mar-2006Daily statusalloncley
33 Thu 09-Mar-2006Daily statusalloncley
32 Thu 09-Mar-2006another TPOP crashbaseoncley
31 Thu 09-Mar-2006TPOP crashbaseoncley
30 Tue 07-Mar-2006Daily statusalloncley
29 Mon 06-Mar-2006school connection weakbaseoncley
28 Mon 06-Mar-2006central powercentraloncley
27 Sun 05-Mar-2006south 30m now serializedsouthoncley
26 Sun 05-Mar-2006strange stuffbaseoncley
25 Sun 05-Mar-2006Serializer WARNING!southoncley
24 Sun 05-Mar-2006sonic S/Nalloncley
23 Sun 05-Mar-2006site visitsalloncley
22 Sun 05-Mar-2006Proceduresalloncley
21 Sat 04-Mar-2006Sonic Updateallagee
20 Wed 01-Mar-2006Daily statusalloncley
19 Wed 01-Mar-2006west visitwestoncley
18 Wed 01-Mar-2006south visitcentraloncley
17 Wed 01-Mar-2006central servicecentraloncley
16 Tue 28-Feb-2006central boom anglescentraloncley
15 Tue 28-Feb-2006usb drive installsouthoncley
14 Tue 28-Feb-2006usb drive restartwestoncley
13 Tue 28-Feb-2006Daily statusalloncley
12 Tue 28-Feb-2006guy wire tensionsallagee
11 Tue 28-Feb-2006network stuff alloncley
10 Mon 27-Feb-2006monitoring toolsbaseoncley
9 Mon 27-Feb-2006Old status messagesalloncley
8 Mon 27-Feb-2006Network configuationalloncley
7 Mon 27-Feb-2006South network tasksouthoncley
6 Mon 27-Feb-2006AP24 settingwestoncley
5 Mon 27-Feb-2006soilsalloncley
4 Mon 27-Feb-2006Daily statusalloncley
3 Mon 27-Feb-2006school internet contactsbaseoncley
2 Thu 23-Feb-2006Phone numbersallmaclean
1 Tue 21-Feb-2006Logbook creatednoneaster

268: data-system, Site west, Mon 01-May-2006 16:09:27 PDT, West data copied
The download of the onboard pocketec data to exthd7 from West was started as of ~1610LT.

Based on data storage levels on exthd7 and exthd9 the data from Central tomorrow may need to be placed on exthd9.
267: data-system, Site south, Mon 01-May-2006 12:42:50 PDT, South data copied
The south pocketec data was copied to /media/exthd7 in the 'south' subdirectory. This is the final download of south data.
266: daily status, Site all, Mon 01-May-2006 08:54:34 PDT, Final status and tear down duties in brief
ISFF T-REX Daily Status 1 May 2006

All systems and sensors are operating normally. 

Teardown begins today with South, West and then Central.

A singly redundant copy of isff... data has been copied to exthd9 and of the final hotfilm run to exthd7 through the end of the UTC day 30 Apr 2006 (the end of T-REX). Discs are quite full but I anticipate sufficient space. The final copies will be made as the pocketecs are gathered today.

I will bring home one copy, and the pocketecs on Friday 5 May 2006.

The Stanford/Berkeley group continues to extract soil moisture/temperature sensors and utilize the ISFF oven/TRIME/scales as the trailer fills. Poulos will finish their gravimetric scales from his hotel room so that they can return for finals, starting Tuesday.
265: HotFilm, Site central, Sun 30-Apr-2006 17:33:22 PDT, Inverter failure ends hot-film run at 1000 LT 30 Apr
The power inverter lent by Oncley to power Ripple in the hot-film/stn0 electronics box was unable to operate due to low battery power of unknown origin. A brief effort to ascertain the genesis of the low power was made approximately 30 minutes prior to the end of T-REX (00UTC 1 May). We soon decided that the benefit was limited given the end of the project (we had originally intended to run until the Central tower was shut down).

The hot-films were gathered by Poulos at ~ 1900 UTC 30 Apr.

Thanks for all the hard work and the gathering of some 300 hours of data.


264: soil, Site all, Sun 30-Apr-2006 11:51:24 PDT, Lent gravimetric/TRIME to Berkeley/Stanford de-install team
This morning we lent the TRIME and gravimetric sampling system to the students soil moisture tear-down team. This team consists of Rica, Megan Bela (applied to NCAR summer intern program), Megan D., Bob Street and Frank Ludwig. Brad Agee joined one of their teams of 3 persons. They plan to de-install those soil moisture sites that during a previous visit have shown to give invalid values. Later, after the EOP ends, they will begin uninstalling working sensors. The remainder will be gathered tomorrow.

They will be utilizing the ISFF oven for the gravimetric samples. In addition, equipment borrowed from Inyo County Water Department by that team will be used, along with larger Inyo County ovens.

The Stanford/Berkeley team has become quite good at utilizing the equipment and was thankful for the help of ISFF throughout.
263: daily status, Site all, Sun 30-Apr-2006 10:36:56 PDT, Daily status
ISFF T-REX Daily status - 30 Apr 2006
Poulos, Knudson, Golubieski, Agee on-site

All sensors and systems are operating normally. T-REX EOP-5 will complete at 1300 LT today and T-REX itself formally ends at 00UTC 1 May.

The OTIHS run is ongoing with the ASU lidar staring at the 3m system for lidar/hot-film/sonic dissipation estimate intercomparisons.
262: power, Site central, Sun 30-Apr-2006 10:26:37 PDT, H2 level
Hydrogen level is at 900psi at 1345 local time on 4/30/06.
261: soil, Site all, Sat 29-Apr-2006 17:08:36 PDT, Trime Measurements, etc.
Took TRIME readings and soil samples at all tower locations.

Central:  1329-1345 local time at site.
0-3cm sample
Total Wt.= 128.1g - container wt. 53.5g
Net Wt. = 82.4g w/o container/rocks/roots
container/tare/wet wt/dry wt(without tare)
e1 1.4,25.1,22.5
e2 1.6,25.0,22.7
e3 1.5,23.2,21.8

3-6cm sample
Total Wt= 136.1g - container wt 26.8g
Net Wt = 108.6g w/o container/rocks/roots
container/tare/wet wt/dry wt(without tare)
d1 1.3,35.9,30.2
d2 1.5,38.5,32.2
d3 1.5,34.7,21.8

Sample #,Moisture%,TDR

South: 1445 - 1507 local time at site
0-3cm sample
Total Wt = 171.5g - Container wt 54.2g
Net Wt = 117.2g w/o container/rocks/roots
container/tare/wet wt/dry wt(without tare)
c1 1.5,35.4,36.9
c2 1.4,36.2,37.6
c3 1.5,40.2,41.5

3-6cm sample
Total Wt. = 118.6g - Container Wt 26.8g
Net Wt = 85.8g w/o container/roots/rocks
container/tare/wet wt/dry wt(without tare)
b1 1.6,28.5,29.8
b2 1.3,28.4,29.7
b3 1.5,27.7,29.0

Sample #,Moisture %,TDR

WEST:  1531 - 1545 Time on site
0-3cm sample
Total Wt = 173.7g - container wt 48.4g
Net Wt. = 108.4g w/o container/rocks/roots
container/tare/wet wt/dry wt(without tare)
a1 1.5,36.9,38.4
a2 1.4,35.0,36.5
a3 1.4,36.1,37.6

3-6cm sample
Total Wt. = 103.3g - Container Wt 29.3g
Net Wt. = 73.9g w/o container/rocks/roots
container/tare/wet wt/dry wt(without tare)
f1 1.5,25.7,26.8
f2 1.5,24.7,25.9
f3 1.5,24.3,24.1

sample #, moisture %, TDR

In oven at 1630 local time

260: daily status, Site west, Sat 29-Apr-2006 17:06:56 PDT, Guy Wire Tensions
Guy wire tensions for West's 30m tower

Inner to outer
SE: 350,290,350,320
SW: 325,250,275,275
N:  340,240,260,280

259: daily status, Site south, Sat 29-Apr-2006 17:04:21 PDT, Guy Wire Tensions
South 30M tower guy wire tensions.

Inner to outer
N:  325,275,250,280
SE: 360,310,350,300
SW: 375,275,340,290

258: LOG, Site central, Sat 29-Apr-2006 15:35:01 PDT, ISFF media day
On late 28 Apr 2006 Poulos and Golubieski took the Haines-Stiles videographer to the Central site. In addition to larger questions regarding the overall success of T-REX an extensive set of high resolution (digital camera quality) video was taken of the Central tower and the OTIHS 5m tower nearby. Detailed pictures of the hot-film anemometers, motors and operations were taken. Kevin said that these were 'cool'. An amusing experience to say the least.
257: daily status, Site all, Sat 29-Apr-2006 15:16:56 PDT, Daily status
ISFF T-REX Daily status 29 Apr 2006
Poulos, Golubieski on-site

All sensors, systems and data plotting routines are functioning well today. The OTIHS run continues today (> 25 hours). EOP-4 has concluded at this time and EOP-5 begins shortly and will run until 1pm 30 April.

Today we'll be taking final gravimetric/TRIME measurements at each tower today.

It appears, after further investigation, that the soil temperature deviation at Central versus West/South (Delta T ~ 10C) may yet be realistic. The soil moisture levels, perhaps due to local groundwater conditions or soil type differences (alkali vs. sand/mixed) seem to be combining to create these differences during daytime. Interestingly, the soil temperature at all three sites converges overnight - implying that the calibration coefficients that have been checked at least once earlier in the experiment, are correct.
256: LOG, Site central, Fri 28-Apr-2006 17:22:32 PDT, Central Guy Tensions
Tensions on 30M tower

NW guys (inner - outer): 400,340,310,300
SE guys (inner - outer): 340,260,250,300
SW guys (inner - outer): 400,320,275,275

Taken approx. 1630 local time.

255: krypton, Site central, Fri 28-Apr-2006 17:19:52 PDT, Krypton Cleaning
30M Krypton clean 1615 LT.

254: daily status, Site all, Fri 28-Apr-2006 15:58:32 PDT, Daily status
ISFF T-REX Apr Daily Status
Poulos, Golubieski on site

All sensors and systems are operating well.  OTIHS run began at ~1230 LT and expected to continue until data system is longer running.
EOP4 began today at 1600 LT and will continue to 1300 4/29.  EOP5 will run 1600 LT through 1300 4/30.

253: krypton, Site all, Thu 27-Apr-2006 18:50:37 PDT, Krypton cleanings - South, Central, West
The krypton hygrometers were cleaned as follows:

Central, 5m (Agee) - 1600-1607
Central, 30m - not cleaned due to high winds

South, 5m and 30 m (Poulos) - 1645-1705

West, 5m and 30 m (Poulos) - 1745-1800

The cleanings seem to have had a significant effect at the south and west tower kryptons(noted as sudden increases in kh2oV).
252: TRH, Site west, Thu 27-Apr-2006 14:41:19 PDT, T/RH 30m West restarted
The T/RH sensor at West 30 m became uni-valued at approximately noon local 26 Apr 2006. With the help of S. Oncley it was master restarted electronically at ~ 11am local 27 Apr (through rserial) to its former level of performance.
251: daily status, Site all, Thu 27-Apr-2006 09:22:01 PDT, ISFF T-REX daily status 27 Apr 2006
ISFF T-REX Daily status 27 Apr 2006

Poulos, Agee on site

All sensors except the 30m T/RH at the west tower are operating normally. Data systems and online data plots are operating well.

IOP-15 ended at 12UTC today.

Krypton voltages appear to be getting low at 30 m for South and West, and at 5 m at Central.

250: data-system, Site all, Wed 26-Apr-2006 16:53:43 PDT, Redundant copies for exthd7, 8, 9, 10
All hard-drives exthd7, 8, 9, 10 have been check_dirs and updated (through the 25 Apr hot-film run and the latest pocketec downloads from each tower).
249: radiation, Site all, Wed 26-Apr-2006 15:11:20 PDT, Radiation Sensor Cleaning
All of the radiation sensors were cleaned today.  South was cleaned at 12:25PST, West at 2:25PST, and Central at 2:43, 2:44, and 2:45 from south to north, respectively.  The net rad (1m) at central was particularly dirty to the point that it was difficult to see through the dome.  Unable to clean the 10m due to being at the site alone.  

248: soil, Site all, Wed 26-Apr-2006 15:09:50 PDT, Soil Sensor Photos
Today I took pictures of the soil sensors at each site.  There are 3 photos for each site, one each from the SW, N, and SE, respectively.
247: soil, Site south, Wed 26-Apr-2006 14:59:36 PDT, Trime Measurements, Gravimetric
25 April, 2006
South Site

Sample taken at 1220PST
Samples placed in oven at 1300PST

0-3 cm  154.8-48.7 = 106.1g net

		Wet	Dry

	D1	30.5	30.2
	D2	33.7	33.5
	D3	41.9	41.5

3-6 cm 143.0-29.6 = 113.4 net

		Wet	Dry

	E1	39.6	38.3
	E2	37.8	36.4
	E3	35.0	33.6

TRIME measurements

SW of soil sensors.

	1 = 2.3
	2 = 2.3
	3 = 2.3

N of soil sensors.

	4 = 3.7
	5 = 3.8
	6 = 3.8

SE of soil sensors.

	7 = 2.9
	8 = 2.9
	9 = 3.0

246: wireless-network, Site base, Wed 26-Apr-2006 12:56:35 PDT, eantE DLR lidar Etherant returned
The DLR lidar tower and etherant (eantE or 14.37) has been returned and is placed under the Base trailer. The antenna will be detached and placed in the back of the base.
245: base, Site base, Wed 26-Apr-2006 11:40:53 PDT, Aster outage when docking station hardware suddenly and permanently failed
At 615am local time it was observed that Aster at the Base trailer was off. The author had never seen Aster shut down the entire experiment and had not shut it down when departing the trailer at approximately 2000 25 Apr.Approximately 2 minutes later the sounding team, Bob Street, Stanford and Serena Chew, DRI, showed up for the morning sounding during the current IOP of T-REX. They noted that lightning had occured overnight, near MAPR, after Midnight local.

An unsuccessful attempt was made to turn on the system at ~0620 after checking power cords and connectors (noting that the power was on to all drives, the monitor, the modems, the Terabeam, the lights, the Leeds sounding computer and various surge protectors. Thus, it appeared that either Aster or the computer port had experienced a catastrophic failure of some kind. Further investigation revealed that
1) the Aster machine would not turn on with battery power (it had drained after power to it had failed),
2) the computer Docking Station would not turn on with when powered from any plug, including known functional power
3) that there was no access to a fuse on the Docking Station through any panels etc.
4) the original 90W power cable for the Aster (Dell Latitude D800) was not in the trailer (two Poulos searches, one Chew search),
5) the power cord for Poulos' Dell Latitude D410 would apply power to Aster, but was of 65W and therefore subject to lesser performance
6) internet access was down in the trailer,
7) Agee reported from online in Bishop that the data plots had been down since before midnight (~2300 local), which when combined with the fact that no other electronic equipment was down, eliminated to a significant degree the 'lightning' theory,
7) #5 allowed some power to the battery and Aster could be functional - leading to the conclusion that indeed the Docking Station had a catastrophic failure.

  After lengthy trouble shooting and Semmer looking up the viability of running Aster on 65W power, we decided to bring up Aster on 65W power. The only peripherals plugged in were the 12.0 and 14.0 connections to limit power usage. 

  Fortunately, this worked after a 2nd hardboot. The GPS time unit was subsequently plugged in as a crucial element of timekeeping after consultation with Gordon. This created no power or performance issues on sub-optimally-powered-Aster.

  While no data had been written to Aster over the wireless, pings, file listings and data_stats of the 3 towers indicated that indeed the local data storage was complete overnight during the Aster outage. No data had been lost and soon thereafter files were filling on Aster with on-line data plots updating.

  Subsequently, it was decided after discussions with Gordon/Steve S. to one-by-one attach various peripherals to Aster in its weak power state in hopes of retaining near normal functionality. An attempt was first made to enable the use of the 19" computer screen rather than the Aster laptop screen - this was unsuccessful - with electronic gibberish being the onscreen result. We thus are using Aster without the big screen until Golubieski arrives with the 90W power supply from Boulder (none were available at in Bishop when Agee arrived first thing this morning and it would be 4-5 days to ship).

  Agee also used a multi-meter to assess the circuits on the Docking station power supply. He found 'beeps' (circuit completion) when the meter was attached to the cord itself plugged in, but not when attached to the male pins at the back of the Docking Station. Our preliminary conclusion is that the internal power supply has failed or a fuse has blown within the Docking Station. Once that occured, apparently randomly, battery power allowed Aster to function temporarily while the battery ran down to zilch. Thus the state at 0615 26 Apr.

  Gordon believes that since exthd7 and 9 are powered separately that their draw will be limited from the USB when plugged. Steve S. has suggested doing this in half-hour stages to ensure Aster functionality at each step.

  We also agreed that since the data acquisition is quite solid onsite, an IOP is underway and the Central tower wireless still potentially shaky after yesterday's tests, that off-loading local storage is not advisable at this time to fill in the online data plot gaps.

244: daily status, Site all, Wed 26-Apr-2006 11:35:04 PDT, Daily status
T-REX ISFF Daily Status 26 Apr 2006

Poulos, Agee on site

All sensors are operating normally at this time.

In the past 24 hours we have 1) executed a series of wireless tests to fix a flaky wireless link to the Central, and 2) experienced a hardware failure of the Aster docking station (at approximately 2300 local, last night). Item #1 resulted in the loss of ~ 1.5 hours of Central data (all) and an additional 4.5 hours of 5m and 15m sonic data at Central mid-day 25 Apr. Item #2 resulted in a large gap in the online data plots until the Aster system could be restarted with a partial power (65W) cord from Poulos' much smaller laptop. Golubieski will be bringing a full 90W AC/DC power cord from Boulder tomorrow and we will return to near full capability.

Agee has been given responsibility for the following field duties today:
1) Triply redundant TRIME and gravimetric measurements at the South tower
2) Pictures of the soil sensor installation area at South, West and Central
3) Checking the H_2 level at Central
4) Checking spotting of the radiation sensors at South, West and Central (excluding the tall rad sensor) from recent rains and cleaning as needed
5) Obtaining replacement batteries for the GPS and a few extras for a supply in the Base.

243: wireless-network, Site central, Tue 25-Apr-2006 18:22:29 PDT, Wireless fiddle tests
Poulos and Agee executed tests where the Central and Hot-film etherants were switched between HotWire and NIDS SSIDs. We also manually altered the elevation and azimuth of the high gain antenna. 

In all cases while occasional improvements to dB were found (as high as 18dB), the system dB always varied significantly and to low levels (down to 6-8dB). No preferential orientation could be systematically established.

It was noted that REAL and West continued to maintain 20dB+ levels.

The DLR antenna has been returned and we could try it on Central since it consistently returned 20dB+ as well, but given the late state of the experiment we will likely be satisfied with reboots of the Central etherant 12.40 until the end.
242: wireless-network, Site central, Tue 25-Apr-2006 15:23:21 PDT, Wireless intermittent outages, return to the original configuration 12.40
Overnight the performance of the recently implemented eantJ 12.42 at the Central tower was flaky. Although initially a steady performing antenna with between 14-16dB, this system gradually reduced in effectiveness and was found at ~10dB this morning.

The plan then was to move this antenna to the high-gain antenna, which, with little hot-film traffic would be relatively free for Central. In addition, the high-gain antenna could increase the SNR of the new Central antenna by as much as 6dB ideally. With John's help Greg implemented new SSID that was incorrect and shutdown the wireless link to Central. We used 'hotfilm' instead of the actual name 'HotWire' and rebooted unsuccessfully.

As a result the central tower was visited and the ToughBook brought up into the ethernet board, which showed that 12.42 could not be pinged. An attempt was made to try a ping from the Central NIDS to no avail. The system was hardbooted in an attempt to recover functionality, unsuccessfully. As a result some data was entirely lost (roughly 45 minutes). 

Data began coming in again after switching the cable to the original antenna, 12.40, eantH, which had been left up on the tower disconnected when eantJ (12.42) was put up, just in case.

We found that 12.40 came right up with activity in all three lights, whereas reconnecting 12.42 there was 2 lights flickering and the wireless activity light not on. This indicates a wireless configuration problem, I suspect.

12.42 was detached and brought back to base. We do not plan to work on reconfiguring it here now since 12.40 is operating with 12-14dB currently.

We have not attempted to use HotWire as SSID with 12.40 yet.

241: daily status, Site all, Tue 25-Apr-2006 10:27:14 PDT, Daily status 26 Apr 2006
All sensors are operating at this time. 

The wireless to Central continues to perform poorly at intervals. We are switching it to the high-gain antenna in an attempt to improve performance further.

It appears that South and West soil temperature may be miscalculated (their values are in a reasonable range but are behaving non-physcially. The coefficients will be checked.
240: HotFilm, Site central, Mon 24-Apr-2006 19:28:23 PDT, Hotfilm run
24Apr06, 17:40PDT, Central
Began hotfilm test..

End. 25 Apr 2006 ~230pm local.

Dual backup of data completed 26 Apr 2006

239: krypton, Site central, Mon 24-Apr-2006 19:27:03 PDT, Cleaned 5m Krypton
24Apr06, 17:30PDT, Central

Cleaned 5m krypton on short tower
238: rain, Site central, Mon 24-Apr-2006 14:56:19 PDT, Rain Check, Download Data
24Apr06, ~13:45PDT, Central

Yesterday we had 
a) good rain 
b) lots of missing data from central (see other logs)
so darn, we couldn't see the rain quantities!

Today we checked the rain gauge.  It had roughly a residual of ~1/3 bucket,
and we did a tip test (offline from NIDS)......OK, working.

So we halted the viper and swapped in the 4GB usb stick to bring in and
merge the local storage from yesterday....more later...
Yup, the events are there.

237: soil, Site central, Mon 24-Apr-2006 14:49:36 PDT, Tsoil Coefficients, etc.
24Apr06, 13:40PDT, Central

Greg has noticed that the Tsoil readings have appeared quite low during the
entire experiment.   Steve Oncley reviewed the readings and agreed.  Discussion
with Steve Semmer suggested the possibility that the coefficients in the logger
could be incorrect, so we checked:

Sensor Installed per label on cable= STP97001, same sensor as spreadsheet
indicates it should be.
Coefficients were exactly as what the 'excel-spreadsheet' said they should be:
Tbl4-Location	Value		Spreadsheet
30		-244.0		-244
31		+235.89		235.89
32		+7.6499		7.65
The spreadsheet file is in/called:
-rw-r--r--  1 aster users 31744 Apr  6 16:00 radiationinfo.xls

Re: wiring
A subsequent check and the wires appeared OK.  I did notice that the E3
excitation seemed a bit odd to me: 4-each 5mS pulses spaced at 5mS apart
and at the correct 2.1VDC.

236: soil, Site central, Mon 24-Apr-2006 14:39:55 PDT, Trime Measurements, Gravimetric
24Apr06, 13:30, Central

~14:15PDT soil sample taken
Definitely moist from rain yesterday, so good opportunity.  Seems quite well
formed and clay-like except for the 'loose alkaline top.'

0-3cm	125.0-45.1 = 79.9gm
	e1	25.8	20.0
	e2	15.8	12.4
	e3	18.5	14.4

3-6cm	169.8-73.1 = 96.7gm
	d1	20.2	14.9
	d2	21.3	15.7
	d3	18.5	13.6

14:00PDT Trime Greg took these readings:
Sample	%Vol	TDR
1	39.8	24
2	42.3	13
3	39.8	0
4	39.6	24
5	42.1	12
6	39.8	0
7	39.6	24
8	42.1	12
9	39.1	0
10	38.2	22
11	38.7	22
12	39.3	22

All 4 sets of 3 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.
#1-3, #4-6, #7-9 were in separate locations.  Because of the funny 0 TDR
values as the measurements proceded, the last set #10-12 were taken.
Ground definitely has the moisture in it.

235: soil, Site west, Mon 24-Apr-2006 12:07:11 PDT, Trime Measurement, Gravimetric
24Apr06, 10:20, West

~10:45PDT soil sample taken
Some pebbles in 0-3cm, nothing in 3-6.  Rain yesterday, so good opportunity.

0-3cm	153.8-53.7 = 100.1gm
	b1	28.0	27.7
	b2	32.0	31.6
	b3	26.6	26.1

3-6cm	136.5-27.0 = 109.5gm
	c1	35.7	34.5
	c2	36.7	35.5
	c3	36.8	36.8

10:20PDT Trime Brad took these readings:
Sample	%Vol	TDR
1	5.6%	
2	5.8
3	5.9
4	4.5
5	4.5
6	4.6
7	5.4
8	5.5
9	5.4

All 3 sets of 3 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.   #1-3 were in bare ground east of sensors,
#4-6 'mixed' west of sensors, #7-9 bare, south of sensors.

234: krypton, Site west, Mon 24-Apr-2006 12:05:55 PDT, 5m, 30m Kryptons cleaned
24Apr06, ~10:25-35PDT

Cleaned 5m,30m with methyl alcohol

233: daily status, Site all, Mon 24-Apr-2006 08:20:05 PDT, Daily status
Daily status 4/23/2006

Poulos, Militzer

All installed sensors are operating normally.
The aerosol sensor at west had its laser go and was returned to the factory.

All data systems are operating.
Wireless outages with Central have improved greatly overnight with (multiple choice)
a) lack of rain
b) slight readjustment of eantj's pointing angle.
c) reboot of tpop
d) coincidence
e) all of the above

232: power, Site central, Sun 23-Apr-2006 19:43:02 PDT, Hydrogen replaced at Central
The H2 at Central was replaced with a full (2200PSI) container. This will last until the last day of the experiment.
231: wireless-network, Site all, Sun 23-Apr-2006 19:08:04 PDT, Rebooted TPOP, Slight adjust of EantJ on central
~17:45 Rebooted TPOP.  Reestablished link with central-eantJ, and sodar-eantg.

~18:30 Realigned very slightly eantJ at central.  Steven Mobbs, watched the
signal strength for us at the time and noted we got roughly 3db improvement.
Also slightly improved the tape job on the connector.  I didn't/couldn't
really pull apart the connector to see if water had gotten inside but my
impression is that it probably did not.

230: wireless-network, Site base, Sun 23-Apr-2006 13:16:29 PDT, changed network route to west
Written from Boulder...

central and west systems have not been reporting on the network for
the last few hours, but south is fine.  Looks like TPOP, which serves is sick again.

west ( is on the AP24 ethernet interface. The base has an 
etherant on both the .12 and .14 networks, so there are two ways to
route traffic to west, either through (AP24 internal
antenna), or (AP24 external antenna).

Up to this point we were routing though, which requires
the TPOP:

$ route
Kernel IP routing table
Destination     Gateway         Genmask         Flags Metric Ref    Use Iface    *        U     0      0        0 eth1    ap24int   UG    0      0        0 eth0    *        U     0      0        0 eth0

I think it would improve things to route through the .14 network and
avoid the TPOP.

Changed the route using these commands:

route delete -net gw
route add -net gw

Also changed /etc/sysconfig/static-routes so that it takes effect on
the next reboot:

any net gw

route command now shows:

$ route
Kernel IP routing table
Destination     Gateway         Genmask         Flags Metric Ref    Use Iface    *        U     0      0        0 eth1    ap24ext   UG    0      0        0 eth1    *        U     0      0        0 eth0

west is now responding.

On west, also changed the address for server in /etc/hosts
from to

In the /usr/local/isff/projects/TREX/ISFF/ops/ops*/ads3.xml, 
changed the destination address to 14.21:


On west, did /etc/init.d/nids restart and it is now sending data.

229: daily status, Site all, Sun 23-Apr-2006 07:20:39 PDT, Daily status
Daily status 4/23/2006

Poulos, Militzer

All installed sensors are operating normally.
The aerosol sensor at west had its laser go and will be returned to the factory.

All data systems are operating.
Wireless outages with Central have worsened substantially.
We installed a different antenna yesterday morning and it ran through the day.
However about 20:00PDT 4/22 it became very intermittent.  Perhaps coincidentally
light drizzle entered the valley at that time with heavier rain this morning.
When conditions allow we'll make another attempt.

228: LOG, Site central, Sat 22-Apr-2006 16:04:43 PDT, Boom angles, central
The boom angles at Central were shot on Saturday morning and Sunday morning (2 and 5 m).

Sonic  Solar   Solar   Targeted     Boom     Corrtn     Boom angle
 Hgt    Time    Azim    Value       Angle    to angl  after correction
------ ------ ------- ------------ -------   -------- ----------------
  1 Sat 1120  134.25d  134d 35.7m  88d 42m    -20.7m   88d 21.3m
  2 Sun 1005  112.11d  112d 27.5m  87d 47.5m  -20.9m   87d 26.6m
  3 Sat 1120  134.25d  134d 35.7m  88d 42m    -20.7m   88d 21.3m
  5 Sun 1025  116.92d  117d 03m    85d 18.5m   -7.8m   85d 10.7m
 10 Sat 1045  122.69d  122d 20.5m  92d 28m    +20.9m   92d 48.9m
 15 Sat 0955  110.21d  110d 44m    92d 59m    -31.4m   92d 27.6m
 20 Sat 1100  127.29d  127d 18m    93d 52.5m   -0.6m   93d 51.9m
 25 Sat 0955  110.21d  110d 44m    92d 59m    -31.4m   92d 27.6m
 30 Sat 1100  127.29d  127d 18m    93d 52.5m   -0.6m   93d 51.9m

227: LOG, Site south, Sat 22-Apr-2006 16:00:35 PDT, Guy wire tensions
With a light breeze from the south, tensions at the south tower were: 

Guy     Orientation
 #     N     SE    SW
 1    330   375   400
 2    275   320   300
 3    290   375   340
 4    280   325   300
226: krypton, Site south, Sat 22-Apr-2006 15:03:35 PDT, 5m, 30m Kryptons cleaned
22Apr06, ~12:00PDT

Cleaned with methyl alcohol

225: wireless-network, Site central, Sat 22-Apr-2006 13:58:33 PDT, Outages continue; New Eant at Central
22Apr06, 8:30-9:30 PDT

The central etherant was down a couple times yesterday and was restarted twice
as late as midnight.  Then it was down again this morning at 5am.
Yesterday I pulled the eantj-sodar off the sodar to swap in at central.

When I climbed to central EantH, here are a couple observations while it
was still 'down.'
- power and wireless activity lights were present, looked good.
- the ethernet connection light was out.
	Note, Greg had wanted to go through swapping the pocketec proceduer
	before doing the antenna, and we had some troubles again
	with getting the swap-in to go.  However the ethernet should have
	been up, esp after a hard reboot of the viper.
	This could indicate a problem perhaps between antenna and
	viper instead of to tpop....maybe...
- the cable end going into the Eant, is improper because there is a gap
	between where the insulation was stripped back and the backshell
	of the connector.
	I taped this connector up speculating that it may be an outside
	possibility that our troubles began after the rain in late Mar,
	with perhaps some moisture getting in?....unlikely.
- 15vdc was OK.
Swapped in EantJ-sodar, installed just below the 'H'.   I pointed it slightly
differently, aiming a bit more down and a bit more to the north.....uncertain
whether this 'adjustment' might help or not.   It may move the pointing more to
the side lobe and perhaps away from interfering sources in town??
Left EantH installed in case it proves to be better than the new one.

Re: SNR seen from the Eants
From briefly monitoring the webpage for EantH yesterday we were saw ~13-15dbSNR,
which is marginal (ie 'yellow').  This afternoon while watching the EantJ,
we've been noticing anywhere from 9-17db reported from the antenna with
nominal around 12.

Re: Using EantJ-sodar instead of EantG-school at central.
I put EantG-school on the sodar yesterday afternoon instead of using it for
central.  The reason is because last summer while doing our range tests, 'G'
had been flakey and had an 'X' penciled on it.  During preparations for TREX
it did work OK in the staging area so decided to ship it hoping it would be
used for 'the other guys' or as a last resort.  Instead it found its way onto
the school, a critical link.   It's not out of the question that some of our
problems with 'School-network' were due to 'G' being flakey??
Meanwhile since yesterday afternoon it seems to be happy at the sodar site
200yds from the base.

224: aerosol, Site west, Fri 21-Apr-2006 15:51:20 PDT, Aerosol Laser Bad, return to factory
The laser for the 9012 particulate counter is bad and needs to be returned
to the factory for repair.

This was determined after discussion with Robert Falbo who is the expert on
the sensor.   He suggested two tests (Note it is safe to work on it because
the laser is self contained.

1) check exhaust flow.

The flow goes through the laser first, through some filters, pump, and flow
meters eventually to exhaust which is by the needle valve / nipple mounted on
the pump.  This needs to be roughly 3liters/min.
This test indicated that the pump is working.
The pump it is driven off a PWM technique which ramps the speed up
or down depending upon feedback from the flow-meter and counter.  This is why
it sounds as if the pump starts up slowly and eventually kicks in.  That had
fooled me into suspecting a clog or bad pump.

2) check current-shunt voltage for the laser.

The laser's health is monitored across TP1 (on the small 'perpendicular' pcb
attached to the end of the laser, and to TP3 (ground, on the main laser
electronics board).  These test points can be identified by the red and black
insulators around them respectively.
The correct voltage should be ~800mV; anything above 1V indicates a bad laser.
With this unit, the voltage was 1.6 dropping to 1.25 as the pump ramped up.
Robert confirmed it is bad and gave us a return authorization #

Return / Contact Information

Met One Instruments Inc.
1600 Washington Blvd.
Grants Pass, Oregon  97526
Attn: Robert Falbo

223: HotFilm, Site central, Fri 21-Apr-2006 14:03:20 PDT, GPS coordinates of the hot-film sonic anemometers
GPS coords under the 1m sonic
222: licor-co2, Site central, Fri 21-Apr-2006 14:01:15 PDT, Pictures around Licor
Poulos took three pictures in the vicinity of the 5m tower near Central where the Licor CO2 is sited ~ 1330pm 21 Apr.
221: LOG, Site central, Fri 21-Apr-2006 11:39:07 PDT, Central pocketec download
Central data from 4/11-4/21 was downloaded to exthd7 on aster on 4/21.

The pocketec from the hot-film system is now at central as local storage.

220: daily status, Site all, Fri 21-Apr-2006 11:30:48 PDT, Daily status
Daily status 4/21/2006

Poulos, Militzer

All sensors are operating normally, with the exception of the West tower aerosol system. Discussions are ongoing between John, Shane and the manufacturer. The hot-film ran successfully overnight and is now shutdown due to increasing southerly flow outside the acceptance cone of the system. The pocketec local storage device from hotfilm was moved to Central to allow data download from Central to Base.

All data systems are operating, although we had an overnight outage of the network connection at Central NIDS at ~1am local 4/21.
This required a hard boot to re-establish at ~ 9am 4/21. No data, except for the reboot process, was lost due to ongoing local storage during the network outage.

Intermittently, we also had outages of the wireless network connection to Central through the flaky etherant there. With consultation to Gordon, we will attempt to change it out with the spare in the ISFF Base (eantG, as 12.40 in place of eantH at Central). 

219: HotFilm, Site central, Thu 20-Apr-2006 19:25:28 PDT, hotfilm running

17:20PDT began new hotfilm run with 'fresh' pocketec disk.
218: krypton, Site central, Thu 20-Apr-2006 19:24:31 PDT, 5m Krypton cleaned
20Apr06,, 17:10PDT

Cleaned with methyl alcohol
217: soil, Site south, Thu 20-Apr-2006 19:16:19 PDT, Trime Measurement, Gravimetric
20Apr06, 18:00-18:20, South

18:0PDT soil sample taken
Some roots in the 0-3cm sample, and pebbles in both.

0-3cm	164.0-79.5=85.5gm
	c1	26.3	26.1 = 0.2gm
	c2	26.5	26.3 = 0.2
	c3	23.6	23.4 = 0.2

3-6cm	160.9-44.9=116gm
	b1	28.6	28.0 = 0.6
	b2	28.6	28.0 = 0.6
	b3	35.2	34.5 = 0.7

11:40PDT Trime:
Sample	%Vol	TDR
1	2.9	64
2	2.9	64
3	6.4	62
4	6.5	62
5	6.9	46
6	7.0	46

All 3 sets of 2 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.
Measurements 1,2 were in a very dry, loose open area.
#2,3 were in an open area under an old dead grass (clipped off to ground),
or and almost 'licen' patch.
#5,6 were taken near and towards a clump of grass under a sage bush.

216: radiation, Site south, Thu 20-Apr-2006 19:13:27 PDT, Rad.cleaned
20Apr06, 17:55

Cleaned all domes.  No real evidence of dirt on tissues or visually on domes.

215: daily status, Site all, Thu 20-Apr-2006 10:56:04 PDT, Daily status
Daily status, ISFF T-REX, 20 Apr 2006
Poulos, Militzer onsite, Maclean left this morning

The aerosol sensor at West has been removed from the tower for checkout in the base station. 

All other sensors are operating normally.
The Hotfilm is continuing to run normally and had very good conditions last night.

214: aerosol, Site west, Wed 19-Apr-2006 20:41:24 PDT, Aerosol down

The aerosol particulate counter has been reporting bad data since
the 4/13.   Values now are almost all zeroes, with some integer values.

Tech support from MetOne sent the manual, which really doesn't say
much, and the windows configuration/monitor software which I installed
on the panasonic.  At the site, this software failed to communicate
with the counter even though procomm was seeing the raw messages.  At
one point I did manage to grab it's help printout during startup:
Tomorrow we'll bring the sensor down to the base

20Apr06, ~9:15PDT
Removed aerosol from tower.
At the base it appears to work, the pump starts slowly, but runs. It could
be that the pump, or filters are clogged because of the slow startup.
Other attempts to communicate with it via 'particomm' software still failed.

213: soil, Site central, Wed 19-Apr-2006 19:41:43 PDT, Trime Measurement, Gravimetric
19Apr06, Central

12:20PDT soil sample taken
Some roots in the sample.

0-3cm	126.0-79.3=46.7gm
	e1	11.6	10.6
	e2	15.2	14.0
	e3	14.9	13.6

3-6cm	151.8-73.1=88.7gm
	d1	20.7	16.1
	d2	16.1	12.5
	d3	20.6	15.9

11:40PDT Trime:
Sample	%Vol	TDR
1	42	16
2	41.7	16
3	44.4	13
4	45.1	13
5	37.2	15
6	36.7	15

All 3 sets of 2 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.
Measurements 5,6 were taken closer to a bunch of grass on the north side.

212: radiation, Site central, Wed 19-Apr-2006 13:33:31 PDT, Rad.cleaned
19Apr06, 12:25-33

Cleaned all domes.  The Rnet and some of the long, short wave radiometers did show a bit of dirt on the tissue.  The Rnet had visible dust on its domes but not heavy.

211: daily status, Site all, Wed 19-Apr-2006 13:00:05 PDT, Daily status
Daily status, ISFF T-REX, 19 Apr 2006
Poulos, Militzer and Maclean onsite

The TP01 west was repaired and returned to functionality as of 6pm, 4/18/06.
The aerosol sensor at West has been down since the afternoon of 4/18/06 - it did not recover after a power cycle.

All other sensors are operating normally (although occasional Etherant outages at Central continue to create gaps in our online plots).

210: soil, Site west, Tue 18-Apr-2006 17:45:59 PDT, Trime Measurement, Gravimetric
18Apr06, West

16:23PDT soil sample taken
Yuck, very dry and lots of gravel.  The 0-3cm had a large 'rock' in it.

0-3cm	169.7-45.1 = 124.6gm
	b1	30.6	30.2 = 0.4
	b2	21.5	21.2 = 0.3
	b3	41.0	40.1 = 0.9

3-6cm	191.7-79.5 = 112.2gm
	c1	29.5	28.6 = 0.9
	c2	17.2	16.8 = 0.4
	c3	32.8	31.8 = 1.0

16:00PDT Trime:
Sample	%Vol	TDR
1	6.9	64
2	6.8	63
3	5.0	63
4	5.0	63
5	9.6	63
6	9.6	63

1,2 didn't have 100% penetration, with ~3/4" of the top exposed.
3,4 was in a more open area, and had the same problem as above except I
    'cheated' and moved a bit of gravel around the top.
5,6 had good penetration and was closer to a bunch of grass, ~8" away
    with the 45-deg 'slant' toward it.

209: radiation, Site west, Tue 18-Apr-2006 17:43:37 PDT, Rad.cleaned, Rnet leveled
18Apr06, 15:50-15:55

Cleaned all domes.  Absolutely no evidence of dirt.

Adjusted level on Rnet.
tilted it 'up' from a sag to the south that had the bubble outside
of the target but not pegged.
208: power, Site west, Tue 18-Apr-2006 17:41:21 PDT, Installed Charge-Monitor, Reboot
18Apr06 15:05-15:15

Installed new charge-control-monitor board to get status values.
Vb,Icharge,Iload,Temp  /ttyS13
Gordon's xml worked perfectly and the plots are now available.

207: sonic, Site west, Tue 18-Apr-2006 15:41:53 PDT, Boom angles shot
Poulos, Militzer, West visit 1600 Local 17 Apr 

We shot boom angles with the new Theodolite.Shots were taken from the road west of the tower at ~ 50 m distance. The original records for the table below are noted in Poulos' green record notebook.

Sonic   Solar    Solar  Solar    Solar       Actual    Resulting   Corrected
Height  time of  Elev   Azim      Azim       Azim      Boom angle  Boom angle
        Record   stor   stor    stor d m     Meas'd    Meas        (cardinal)
------  ------   -----  ------  ----------  ---------- ----------  ----------
05 m     1730    67.15  266.38  254d 46.8m  266d 30.5m  92d 21.2m   92d 13.7m
10 m     1715    64.15  264.02  271d 49.8m  264d 13m    91d 11.2m   91d  0.2m
15 m     1655    60.19  260.74  254d 46.8m  260d 40m    90d 26.5m   90d 31m
20 m     1615    52.38  253.54  267d 19.2m  253d 47.5m  91d 48.5m   91d 33m
25 m     1655    60.19  260.74  254d 46.8m  260d 40m    90d 26.5m   90d 31m
30 m     1615    52.38  253.54  267d 19.2m  253  47.5m  91d 48.5m   91d 33m

206: soil, Site west, Tue 18-Apr-2006 15:10:24 PDT, TP01 wiring, possible fix
18Apr06, 15:20-15:45

TP01 wiring harness swapped in logger.
The 'existing' harness had a problem with the +10V regulator output
needed for the heater.  I confirmed that the spare worked in the lab by
tying the '12Sw.Control' line too +5 and observed the 10V on pin-8.  At
west the same test failed.
Speculation is that the outage since 31Mar06 may be fixed.
It's possible that when the Xbee radio was installed in place of the Xstream
that the flimsy regulator wiring in this harness may have broken or
become intermittent.  We'll see....

Poulos, Militzer - West site visit 1600 LT Apr 17
Last night: 17Apr06
I beeped the wiring to confirm continuity as shown below and tried checking
the output of the regulator but that proved very strange after tying the
supply into +12: I could have sworn I saw 3.5V on Analog Ground....eeek!
The wiring of TP01 inside the logger is:
CR10	Color	Bulkhead Amp-9
4H	White		1
4L	Green		4
AG	Black/Bare	7,9
5H	Brown		3
5L	Black		2
Sw.+12	'Brown'		regulator inside heat shrink, +10 to pin-8
Gnd	Black		7,9

CR10 Other wiring:
C7 = Switched 12 Control

205: LOG, Site west, Tue 18-Apr-2006 15:04:35 PDT, West guy wire tension check & re-tension
Militzer and Poulos, West site visit, 1600 LT, Mon 17 Apr

Guy-wire tensions
1 375
2 320
3 360
4 310

1 340
2 295
3 290
4 295

1 340
2 280
3 280 (retensioned to 300)
4 280 (retensioned to 300)

Note on North 3/4. They were retensioned to ~ 325 and then backed off after
some minor bowing was noted.
204: HotFilm, Site central, Tue 18-Apr-2006 09:54:16 PDT, hotfilm system clock ahead
ntp on daisy was mis-configured, and as a result its
clock was  591.58 seconds ahead.

This was corrected at 09:19 this morning, and the
system rebooted, so data recorded after this time are OK.

I believe it has been mis-configured since it was
put on the 192.168.15 network early in the project.

203: daily status, Site all, Tue 18-Apr-2006 09:30:26 PDT, Daily status
ISFF T-REX Daily status for April 18, 2006

One sensor is currently inoperative: the volumetric soil heat capacity probe at the West tower. All other sensors are operating normally.

A hot-film run began at 3pm LT April 17 and is ongoing (expected to continue through EOP-3 [4pm LT 4/18 - 1pm LT 4/19]) until, perhaps as late as Friday 4/21.
202: daily status, Site all, Mon 17-Apr-2006 05:51:48 PDT, Daily status
Daily status April 16, 2006, with Maclean, Militzer, Poulos on site.

All sensors at all tower sites and data systems are operational at this time. 

See 15 minute updates at

OTIHS (hot-films) were run last night from ~19:00PDT 16 Apr to ~0830PDT 17 Apr. Data acquisition was successful. OTIHS will be redeployed later today as synoptic flow subsides.

Winds during the night rotated the antenna from the base to the hotfilm site, losing the remote link to the system (all data was recorded locally). Velocities remained below 10m/s at the ground level but increased significantly this morning such that OTIHS was shut down at 830am LT. 

201: HotFilm, Site central, Sun 16-Apr-2006 19:32:39 PDT, Started hotfilm
Started hotfilm sampling at 19:30 PDT, Apr 16.

This run lasted until 20Apr06 ~12:30 when it was taken down for awhile to
allow the full pocketec to be off-loaded at the base
200: daily status, Site all, Sun 16-Apr-2006 10:09:50 PDT, Daily status
Daily status April 16, 2006, with Maclean, Militzer, Poulos on site.

All sensors at all tower sites and data systems are operational at this time. Online plots are now updated every 15 minutes at:

IOP #13 is underway and scheduled through today and until 08UTC Monday.

There was an ~2.5 hour wireless outage to Central in the early morning hours and a brief dsm-server outage to all three sites also in the early morning hours. No data was lost due to local storage at the tower sites but the online plots and redundant data storage on Aster show these gaps.

Another OTIHS (hot-film) data acquisition test began this morning in a new configuration where data files are written in 30 minute increments (each file is ~ 250Mb). Recent data acquisition driver changes have stabilized data acquisition. The next run with the hot-film sensors attached will begin Monday without automated data system restarts.

198: sonic, Site south, Sat 15-Apr-2006 19:09:54 PDT, Boom angles
Poulos and Militzer shot boom angles with the new Theodolite. The leveling bubble was noted as slightly misaligned but tolerably so. John will pursue adjustment tomorrow. Shots were taken from the road west of the tower at ~ 60 m distance. Leveling was completed using a method of dual confirmation of boom parallel siting, theodolite leveling and predicted and actual sun targeting reading. The original records for the table below are noted in Poulos' green record notebook.

Sonic   Solar    Solar  Solar    Solar      Actual    Resulting   Corrected
Height  time of  Elev   Azim      Azim      Azim      Boom angle  Boom angle
        Record   stor   stor    stor d m    Meas'd    Meas        (cardinal)
------  ------   -----  ------  ----------  --------  ----------  ----------
05 m     1625    54.75  254.78  254d 46.8m  255d 05m   91d 11m    90d 52.8m
10 m     1810    75.57  271.83  271d 49.8m  272d 00m   90d 02m    89d 49.8m
15 m     1625    54.75  254.78  254d 46.8m  255d 05m   91d 11m    90d 52.8m
20 m     1740    69.56  267.32  267d 19.2m  267m 37m   91d 37m    91d 19.2m
25 m     1625    54.75  254.78  254d 46.8m  255d 05m   91d 11m    90d 52.8m
30 m     1740    69.56  267.32  267d 19.2m  267m 37m   91d 37m    91d 19.2m
197: HotFilm, Site central, Sat 15-Apr-2006 17:32:25 PDT, hotfilm power
ripple (hotfilm Dell notebook) died this afternoon

It's internal battery was discharged. The inverter that provides
power was giving a little beep every 5 seconds or so and the green
led on the inverter would brighten and then dim.

Looks like the voltage at the 5m tower is not high enough to drive
the inverter.  The two batteries that were at the tower were moved
to the 30m tower this morning.  I moved one back over to the 5m
tower and then the inverter was happy.  However, then the main
power breaker on the enclosure would flip. ripple's internal
battery was discharged, so it was pulling it's highest load.

Installed a power 'Y' so that the power to the inverter is coming
straight from the battery, rather then via the aux plug on the
enclosure.   Seems to work. It is running right now.

196: data-system, Site central, Sat 15-Apr-2006 12:59:55 PDT, Short local storage outage
15Apr06, 11:15-12:06PDT

When I reconfigured power at central and booted the DSM back up, the
pocketec drive failed start up properly.
At 12:05 I cycled power on the pocketec and it started up again.

195: power, Site central, Sat 15-Apr-2006 12:47:08 PDT, Reconfigured power to hotfilm tower.
15Apr06, ~11:00-11:20

Took down ripple and the hotfilm system, plus central tower to
reconfigure the power distribution so that both the main tower and the
hotfilm components are being operated through the 'Iload' circuit.
This way we can monitor our 'full' consumption.

At ~13.4vdc the current peaked at roughly 10.8A when the hotfilm
system was booted up.  After that it has settled down to roughly 6.5A.
So we can expect perhaps 7-8A when the voltage drops.  I checked the
pam_reference document and confirmed the trace is rated 10A and
installed a big fuse so it won't blow.  For overnight at least I
unplugged the beacon.  It seems that we don't really need it since the
site is clost to the much taller cell tower at this location, plus FAA
requirements are 200'.

The ripple/inverter are now running off 'Aux1' on the hotfilm power
distribution panel after the fuse was replaced.  Some care will be
needed when booting up the hotfilm box.  Probably best to:
1) unplug Daisy inside box.
2) unplug Ripple/inverter from Aux1
3) turn on box and observe motors initialize
4) boot Ripple
5) boot Daisy

194: radiation, Site central, Sat 15-Apr-2006 12:44:09 PDT, 10m Rad cleaned, photos
15Apr06, 11:00PDT
Cleaned 10m Rad.  No visible evidence of dust.
A full set of 360, etc photos were taken from the 10m perspective.

193: TRH, Site central, Sat 15-Apr-2006 12:38:43 PDT, 5m TRH filter check
15Apr06, 10:30PDT
The 5m TRH was off for a few minutes while inspecting its filter.
Good news, it appeared very clean although there was a bit of this fine dust on the support piece.
If this site's TRH is clean then almost certainly the rest are also.
192: daily status, Site all, Sat 15-Apr-2006 07:09:44 PDT, Daily status
Daily status April 15, 2006,

On site: Maclean, Militzer, Poulos

5m CSAT at West has been down since 9:45PDT 14Apr.  Restarted at 7:30PDT today.
Other Sensors and data systems operating normally at this time.
IOP #13 begins at 13:00PDT and scheduled to last 36 hours.

Another HotFilm test began last night to check edits Gordon made to the a/d code.

NOTE: ISFF ONLINE DATA PLOTS plots are available.
We are increasing the real-time update rate from 1-hour to 15-minutes.

ISFF on-line logbook is available at:
191: HotFilm, Site central, Fri 14-Apr-2006 15:29:32 PDT, hotfilm data system test
Another hot-film data system test was started at ~ 1420pm (without hot-films themselves) during the latest Central visit.
190: power, Site central, Fri 14-Apr-2006 15:16:23 PDT, Central 2m sonic, power usage
John/Greg visited Central at about 1230 and replaced the 2m sonic cable and confirmed working status at ~ 1425pm.
During this visit we ran a series of checks on power consumption and found the following:
The power adapter for Ripple in the Daisy box draws 0.3A
Ripple, with the screen active and the power adapter, draws ~ 1.5A but somewhat more during certain activities.
The total system draws about 1.5A pulsing to 2+ A with disk writes (includes Ripple, Daisy, hot-film motors, etherant, CTAs and the Air-adapter).

All of Central was found to draw ~6A during the day without the beacon. The 35W fuel cell continued to provide 2A. The Lind Auto-air Adapter was found to be making beeps, which we'll investigate.
189: base, Site , Fri 14-Apr-2006 12:27:23 PDT, Hacker Attack - Changed ssh daemon allows
Check /var/log/messages 

I just happened to be here, and saw continuous hack attacks every 5sec of the following type...

Apr 14 11:51:07 localhost sshd(pam_unix)[3286]: authentication failure; logname= uid=0 euid=0 tty=ssh ruser=  user=sshd
Apr 14 11:51:12 localhost sshd(pam_unix)[3288]: check pass; user unknown
Apr 14 11:51:12 localhost sshd(pam_unix)[3288]: authentication failure; logname= uid=0 euid=0 tty=ssh ruser=

Gordon (Mammothing) suggested shutting down the ssh daemon:
/etc/init.d/sshd stop
a man sshd_config
didn't reveal to me at least a specific 'allow-hosts' so instead i added a line to the /etc/ssh/sshd_config file:
	#Editing 14Apr06, jwm, to restrict logins due to observed hack-attack.
	#hopefully this will do the trick without bumping boulder, and locally on site.
	AllowUsers aster@192.168.* aster@128.117.80*

and saved an original copy in:
and finally:

/etc/init.d/sshd start

188: data-system, Site west, Fri 14-Apr-2006 10:51:58 PDT, Local storage retrieved
Measured electric fence battery = 12.66
Retrieved local storage on pocketec for archiving at base.  Swapped in clean drive.
Grabbed soil sample for Boulder tests.
187: daily status, Site all, Fri 14-Apr-2006 08:43:57 PDT, Daily status
Daily status April 14, 2006,

On site: Maclean, Militzer

Sensors and data systems operating normally at this time.
We've noticed that the etherant on central has remained on-line for several days despite frequent 'halts' roughly a week ago.  Factors could be a) luck, b) improved battery/energy situation c) we removed the etherant at the school d) unknown interfering source has gone away...most likely 'a'.

Imitation HotFilm test began last night to test some edits Gordon made to the a/d code.  Halted this morning due to full usbstick and to save power.

186: HotFilm, Site central, Thu 13-Apr-2006 18:05:34 PDT, hotfilm data system test
Hotfilm data system is up - but no probes are installed.

The prometheus is running a modified version 5.91 of the DSCUD 
driver and library.

The old version was 5.9.  I doubt 5.91 has any bug fixes - most
likely it just includes support for a new board from Diamond.

I've made a slight modification to the 5.91 driver - declaring
as volatile two variables that are shared between the 
interrupt routine and the user read method.  The A2D stayed
up all last night and today but we didn't always have sonics
connected, and never had motors connected.

14Apr06 6:45PDT  Ripple halted
Gordon had to use a 4gb usbstick instead of the pocketec for this test.  Since the ingest was down due to full disk, I halted ripple to save power since it's cloudy and the batteries are down to 12.15
185: sonic, Site central, Thu 13-Apr-2006 18:00:30 PDT, 2m Sonic down
Cable for 2m sonic at central failed when the hotfilm system
was being re-installed. The 2m sonic is now only connected to
184: soil, Site south, Thu 13-Apr-2006 13:16:42 PDT, Updates on Echo/Qsoil negative at south
These notes will cover a few days...

Now that we can plot the data more clearly, the soil moisture at south has been reporting negative values the entire project.
Setup spare logger and 'south' soil with spare Echo in base.  This showed roughly 350+ mV in extremely dry soil.   When I tried the probe in soil sample from central it rose to ~1V and consistent with what that station shows.
At south I tried the spare probe with the station logger with the data scope again.  This time I was more prepared to record the signal properly which was not the case yesterday due to wind/connection issues.  Today both probes showed about 380mV or roughly 2.5% per the 'standard' calibration.   I also reproduced this with the spare logger.   I then dumped some water in the 'sample soil' and watched the spare probe go up to ~700mV roughly = 16%, but this seems very low for what was actually in the box.   I also reproduced this off-line with the station logger.

CONCLUSION: my guess is the station Echo probe / logger are working fine.  The soil there appears to be different perhaps because of excess sand or salts (or less than at central).  As a result the standard calibration is probably not correct (refer to Decagon's literature which mentions this issue) don't believe the reported values!
We will bring sample containers of representative dirt from all 3 sites back to Boulder for later evaluation.

Checked Echo/logger at south
E2 excitation - per data scope is getting up to about 2500mV, and reading is at ~50mV.  Appears to confirm the logger excitation is going high enough.
Soil Sample - taken at both south and central.  Central was definitely much more moist, but to the untrained eye the soil itself appeared similar.
Spare Echo probe comparison - The signal from both the installed sensor and the spare inserted in the 'sample soil bucket' were the same: roughly 25mV raw and -40% per the logger message.  This reported value seemed a bit low per the above, so will try testing in the base tomorrow.

Rough comparison of Echo Probes between south and central per meter.
Wiring at south was OK, although the input wire on logger 11H channel was a bit loose and needed a slight tightening.
South:		139mV at highest
Central:	300-350mV at highest
Excitation is supposed to be 2500mV and the readings roughly 10-40% of that per the manual.  The slow meter was unable to grab a proper 'max' of the excitation voltage on E2 of the logger.

182: power, Site central, Thu 13-Apr-2006 12:36:21 PDT, Update on 35W Central Fuel Cell
Central 13Apr06, 10:30 PDT
Last night we said...
(after putting 35W on main batt bank) We'll observe overnight and expect some charging for central out of this unit.  If it works, we may move the fuel cell and batteries adjacent to the rest of the stuff and use the long cable when needed to run the hotfilm system.

Overnight it did provide 1A charging despite the long cable and blocking diode.
Today we moved it to the main battery bank, and removed the blocking diode going into the Icharge2 circuit.

Currently the config. is the solar is charging through the prostar on Icharge1,
and 35W f.c. charging through the '8A' charger in it's box, on Icharge2

The fuel cell then began charging at ~2A, and the solar at ~13.5A
swapped tank in:
2500psi, regulator set at 15psi to f.c.
181: sonic, Site central, Thu 13-Apr-2006 12:34:38 PDT, Guy Wires added to 5m tower
13Apr06, ~10:30PDT

Installed 2 'cross' guy wires to 5m tower to try to help stablizing the oscillations in the wind.

NOTE: this may have effected the sonic levels as the cables were tightened.
180: krypton, Site central, Thu 13-Apr-2006 12:19:17 PDT, 30m krypton cleaned
13Apr06, ~10:00PDT

- 30m Krypton cleaned.  Voltage went from ~.9 to ~1.1 vdc

179: daily status, Site all, Thu 13-Apr-2006 12:17:55 PDT, Daily status
Daily status April 13, 2006,

On site: Maclean, Militzer

Sensors operating normally at this time.

Added extra guy wires to the 5m tower @ central to
reduce the twist oscillations.

60W fuel cell in base trailer for testing, shipping to manufacturer.

Hotfilm data system in base for testing.
178: power, Site central, Wed 12-Apr-2006 11:37:10 PDT, Update on Central Fuel Cell / Battery Status
Central 12Apr06, 9:00-14:30 PDT
Last night we said...
"For overnight we'll run the fuel cell straight onto the battery and hope it comes to life."

It never really did get going and at most provided .5A charging.  This morning the tank has about 2400psi.  We fiddled with the controls, pressures, etc. some more but decided this '60W' fuel cell requires more effort than is advisable in the field plus since it isn't doing anything, and given the note several have made about it's 2 'questionable' internal cells, we removed it from central and brought it back to the base.

35W fuel cell put on central, ~9:50PDT
Since Greg reports no hotfilm runs are imminent, we reconfigured to put the 'happy' 35W f.c. through the central charge monitor 'Icharge2' circuit in place of the bad unit.  This is going through a long (~100') cable, and the blocking diode and then straight to the battery at the end of the chain.  We still have the 2 batteries tied directly on at the 35W f.c.   Given the voltage drop of the cable and the diode, plus the solar charging in full sun we are seeing:
.85A charging from 35-f.c.
13.4 VDC at batteries on f.c.
12.7 VDC at battery on Icharge2
1100 psi remaining in tank.

We'll observe overnight and expect some charging for central out of this unit.  If it works, we may move the fuel cell and batteries adjacent to the rest of the stuff and use the long cable when needed to run the hotfilm system.
177: radiation, Site central, Wed 12-Apr-2006 11:33:03 PDT, 'cleaned' 10m K&Z
12Apr06, ~10:15
attempted to clean the radiometer with methanol using Steve Semmer's sponge mop.
The wind was substantial so I only got the downward looking dome swiped.  There was no evidence of any dirt either by visual inspection or by looking at the 'swabber.'  Consequently I did not maneauver around the logger, etc. to swipe the upper looking dome.   The suspicion is there is just too much wind for this very fine dust to stick, and not enough moisture.   We have not seen any rain.
176: krypton, Site central, Wed 12-Apr-2006 11:31:02 PDT, 5m krypton cleaned
12Apr06, ~9:40 PDT
cleaned 5m krypton on short tower with methanol.
Voltage changed from ~.2 to ~.5, not spectacular.

Note: very windy and this short tower swaps a great deal.

175: daily status, Site all, Wed 12-Apr-2006 08:23:24 PDT, Daily status
Daily status April 12, 2006,

On site: Maclean, Militzer

Most sensors operating normally at this time. The 5m Csat at Central was down overnight from roughly 17:00-6:30.  Data messages were coming through, but all flags were set.  The 5m krypton running off this sensor shows a large drop in voltage (.4) at this time but the trace otherwise is believable. 

The blankety-blank Fuel Cell at Central provided a whopping .5 A charging current at best overnight.

174: power, Site central, Tue 11-Apr-2006 15:41:20 PDT, Update on Central Fuel Cell / Battery Status
Notes for 2-days regarding the fuel cell:

Central 10Apr06, 14:00-14:30 PDT
Installed New charge controller with SOAD onto /ttyS13
Swapped in new H2 bottle: 2500psi, regulated at 20psi
Current from f.c. no more than .15A

Note: we did observe a very brief time where the fuel cell 'claimed' to have provided about 4Amps charging, but only for a few minutes.  After that it stopped.

It was very very nice to finally have the SOAD reporting Iload, vbatt, Icharge1 (pv) and Icharge2 (cell).....thanks Gordon!

11Apr06, 13-14PDT
H2 bottle down to 900psi.  No charging overnight. VOC=~16
Is it simply purging the bottles??

In theory that going through prostar charge controller may be a problem, we installed a segregated set of wires through Icharge2 directly to the end of the battery chain, avoiding the controller entirely.  We also swapped in a new H2 bottle.  After restarting we observed on both the SOAD and directly on a meter ~.45A charging to the battery.

NOTE: the cell-cell voltages showed most at ~.75v, but one was at -.15 and another was fluctuating from ~.2-.75.
Possible that these 'bad' cells are draining the current exactly like a bad battery cell would do.  Need to measure their impedances.

NOTE: We did observe a slight, but noticable increase of roughly .15+ A out when we openned the lid of the electronics box enclosure.

For overnight we'll run the fuel cell straight onto the battery and hope it comes to life.

173: wireless-network, Site base, Tue 11-Apr-2006 11:31:58 PDT, Removed network equipment from school
John and I removed all the network equipment from the Independence
school yesterday.  That link was not being used and it gives us
a spare etherant WIFI antenna. 

172: daily status, Site all, Tue 11-Apr-2006 09:24:56 PDT, Daily status
Daily status April 11, 2006,

On site: Maclean, Militzer

Most sensors operating normally at this time. The 15m Csat at Central was down until this morning after we power cycled the station at ~14:00PDT yesterday afternoon when we installed the power monitor board. 

Hotfilm run began last night at ~19:30PDT (2:30Z today).  We will remove it later this morning.

171: HotFilm, Site central, Mon 10-Apr-2006 19:59:30 PDT, Hotfilm run
Begin 10Apr06,19:45 PDT Started hotfilm.
End   11Apr06 12:15 PDT

The fuel cell running ripple is looking good: 1200psi remaining, and the 2 batteries are at 13.6 on the f.c. before we started.  At the end, the batteries were still 13.6 and the f.c. 1100psi.
The main system batteries were at 12.55, and thus ~11.55 going into daisy at the end of the long cable when we began.

170: data-system, Site all, Mon 10-Apr-2006 11:59:51 PDT, restarted data systems
Towers and base are now running new NIDS code

New code has the StatisticsProcessor and support for
prompting serial sensors (needed for SOAD charge monitor).

169: daily status, Site all, Mon 10-Apr-2006 09:19:38 PDT, Daily status
Daily status April 10, 2006,

On site: Maclean, Militzer

All sensors operating normally at this time.  Wireless link at central went down yesterday afternoon again and required a power reset.  As usual, there was no data loss on the local storage.

168: power, Site central, Sun 09-Apr-2006 18:36:56 PDT, Update on Central Fuel Cell / Battery Status
The Batteries at Central have been gradually charging with the full sun and the hotfilm system off most of the time.

Fuel Cell System at Central
Note: this note begins from 7Apr when I did a check after dark.

19:40 PDT, after dark.
	Fuel Cell VOC     = 14.4	seems OK
	Fuel Cell Icharge = 0		bad. despite batteries at 12.5
	Metered open ground on Icharge2
	Moved Fuel Cell to 'solar-in'
	Fuel Cell Icharge = .12A	yuck. thus the non-consumption of H2
					the last several days.
	0) Central has been running only on solar.  The solar system here is
		just enough based on manual battery monitoring the last few
		days.  The added load of daisy/ethant/hotfilm puts it on the
	1) The Fuel Cell is basically only keeping it's fans running at this
		point.  The filter masks are in place but noticably dirty.
		It's possible if I'd left the f.c. connected to the solar-in
		port longer that it may have 'started-in' but that's T.B.D.
	2) The charge monitor board isn't wired properly to allow the F.C.
		to work even if it were running ok.

	The next day, Apr8th, we removed the charge monitor that has an SOAD
	from the fuel cell box and brought it to the base to get running.  After
	making a cable for the SOAD - NIDS we confirmed the serial 
	interface/command and Gordon began work on 'polling software' for NIDS.
	This will allow us to use the SOAD instead of fudging in a logger for
	the power monitor, plus gives another important NIDS/EVE software
	capability.  By Sunday evening Gordon had this code working on the
	test box.
	We 'fixed' the charge-monitor board so the 2nd charging supply (f.c.)
	can work and added the blocking diode.  Intend to install tomorrow.

167: LOG, Site all, Sun 09-Apr-2006 17:48:03 PDT, guy wire tensions
Measured guy wire tensions yesterday/today, April 8-9

West	   From,March30th	Apr9th, west breeze
	se	ne	w	se	n	sw
1	395	350	360	380	375	350
2	330	285	300	275	280	300
3	395	300	290	350	375	300
4	350	300	300	300	300	300

Central	   From,March30th	Apr8th, lt south breeze
	se	n	sw	se	n	sw
1	430	460	435	410	460	440
2	330	320	350	320	300	350
3	300	345	315	310	340	310
4	305	295	310	340	300	330

South	   From,March30th	Apr9th, 
	se	n	sw	se	n	sw
1	400	355	420	400	340	390
2	350	300	300	340	300	300
3	400	280	350	390	290	340
4	340	300	310	325	310	300

166: radiation, Site west, Sun 09-Apr-2006 17:45:35 PDT, Radiometers 'Cleaned'
16:30PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.

165: radiation, Site central, Sun 09-Apr-2006 14:33:50 PDT, Radiometers 'Cleaned'
11:30PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.
Note: Did not clean 10m rad

164: radiation, Site south, Sun 09-Apr-2006 12:42:28 PDT, Radiometers 'Cleaned'
10:00PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.
Note: slightly adjusted Rnet level
163: soil, Site south, Sun 09-Apr-2006 12:35:29 PDT, gravimetric soil moisture
Took soil sample at south at ~10:20 PDT, April 9
Taken in 'bare soil' similar to where probes appear to be.  There were some small pebbles in the batch; removed before weigh-n-bake

0-3 cm
Total weight = 157.1 - 45.0 gm = 112.1 gm
	wet	dry
d1	26.5	26.0
d2	26.0	25.5
d3	31.1	30.5

3-6 cm
Total weight = 182.2 - 79.5 = 102.7 gm
	wet	dry
c1	25.7	24.8
c2	31.0	29.9
c3	26.7	25.9

162: daily status, Site all, Sun 09-Apr-2006 09:18:45 PDT, Daily status
Daily status April 9, 2006

On site: Maclean, Militzer

All sensors operating normally at this time.  Wireless link had been going down but remained up through the night.  Battery voltage have risen in the last few days with solar charging, may be a factor.

Hot film run began around 1700 PDT yesterday and ran successfully through
the night with some drop outs in the morning.

161: HotFilm, Site central, Sun 09-Apr-2006 09:05:57 PDT, HotFilm Run Sa. 4/8/06 - Su. 4/9/06
4/8/06 ~17:00PDT Begin
4/9/06 ~11:30PDT End of run

Fuel Cell began at 1400psi for ripple supply.
Fuel Cell end   at 1250psi, Vbatt for 2 running ripple=13.5 on F.C.
Central power system for Daisy, etc. was looking good also after 3 sunny days of solar charging.

160: soil, Site central, Fri 07-Apr-2006 18:21:23 PDT, gravimetric soil moisture
Took soil sample at central at ~15:50 PDT, April 7

I intentionally took this sample near where the probes enter the ground and a sage bush.  There were dry stems and roots in this sample, esp in the 0-3 level.  These were for the most part removed for the 'shake-n-bake'

0-3 cm
Total weight = 70.3 - 42.9 gm = 27.4 gm
	wet	dry
b1	6.2	5.4
b2	8.5	7.4
b3	11.5	10.1
Note: I think this 0-3cm is bogus because: 1) procedural error on my part, 2) very dry soil and high winds blew it while transferring.   I.e. don't believe the bulk because of 'suspicious volume'.... this one might be for practise only.

3-6 cm
Total weight = 177.1 - 73.0 = 104.1 gm
	wet	dry
e1	26.0	20.8
e2	21.8	17.4
e3	20.0	15.9
Note: This one was better and should be OK.  It did have some roots and a small worm that needed rescuing before cooking the batch.

159: power, Site central, Fri 07-Apr-2006 17:15:48 PDT, Central Battery Status
Battery voltage checks at Central today:
We have had full sunshine yesterday and today so the charging system should be running full tilt.
~ 11:30 PDT  VDC = 12.60 at 'end' battery
~ 15:30 PDT  VDC = 12.66 at 'end' battery
~ 18:00 PDT  VDC = 12.95 at battery adjacent charger
	Sun going down in sky.
	Icharge-Solar = 7.5A going in
	Icharge-F.C.  = 0.0A
	500psi on F.C. in am.
The end battery is running the daisy/motors so has had a bit higher load the last few days.

~ 19:40 PDT  VDC = 12.45  after Shutdown Daisy/Ripple

158: HotFilm, Site central, Fri 07-Apr-2006 17:09:54 PDT, Imitation Hotfilm Test Running
Winds picking up today.
Begin ~22:30Z

No hotfilms installed and CTA's not running but otherwise daisy/motors/ripple are up so the Gordon can install some new scripts for improved data recovery over the periodic reboot method.

Note: Fuel cell was still running OK.  By that time the batteries were at 12.7 on the f.c. charging.   Swapped in 2 fully charged batteries.

157: HotFilm, Site central, Fri 07-Apr-2006 17:01:08 PDT, Hotfilm Shutdown
Winds picking up today.
Stop Test: Apr 7th, 097: 18:35Z
Started: Apr 5th, 

Note: Fuel Cell was barely running when I arrived.  The regulator showed 0-psi on both gauges.  The battery voltages were at 12.2 at that point.  When I checked the valve it was barely cracked open.  When I openned it all the way the pressure went to 1600psi and etc.  At that point I had shutdown everything.

156: soil, Site west, Thu 06-Apr-2006 15:33:40 PDT, gravimetric soil moisture
Took soil sample at west at ~1500 PDT, April 6

0-3 cm
Total weight = 147.5 - 42.9 gm = 104.6 gm
	wet	dry
d1	25.4	24.8
d2	23.0	22.4
d3	21.4	20.8

3-6 cm
Total weight = 190.5 - 73.0 = 117.5 gm
	wet	dry
c1	21.8	20.4
c2	21.5	20.2
c3	24.3	22.5

155: rain, Site central, Thu 06-Apr-2006 10:55:28 PDT, no rain data at central
The rain gauge at central has been connected to the wrong channel on the 
campbell data logger.  This was fixed this morning.

154: wireless-network, Site central, Thu 06-Apr-2006 10:45:41 PDT, wireless network from central down
The wireless network from central was down from about 9:25 - 10:35 PDT.
Steve and John restarted by cycling power on the etherant at central.
153: daily status, Site all, Thu 06-Apr-2006 09:24:43 PDT, Daily status
Daily status April 6, 2006

On site: Horst, Semmer, Militzer

All sensors operating normally at this time.  Wireless link to central went downyesterday.

Hot film run began around 1800 PDT yesterday and ran successfully through
the night.  

152: HotFilm, Site central, Wed 05-Apr-2006 18:34:19 PDT, Hotfilm running
The hotfilm is up and running at ~18:00.

151: wireless-network, Site central, Wed 05-Apr-2006 18:02:14 PDT, central communications lost
Wireless communication to central lost around 9 am PDT, restored around
1800 PDT.
150: radiation, Site central, Tue 04-Apr-2006 16:09:33 PDT, ZigBee now power on/off at Central
The two single pygreometer logger boxes, #1 and #2, now have their radios
being turned on then off by the logger. They are setup to be on for 1 second
and off for 4 seconds. New batteries were insatlled in each unit. The
code change can be found on the Panasonic under the PC208W directory.
The file name is TREXzig.

149: soil, Site west, Tue 04-Apr-2006 11:14:57 PDT, missing data from logger at west
I am investigating the missing logger data reported on the Splus plots.
It appears that the data archive is complete, while the Splus file is
missing some data.

(P.S. This problem appears to be "fixed" by re-running stat_proc.)

Here is data from west logger as seen by Splus

               Time(PDT)     Tsoil    Qsoil
 04/04/2006 04:57:30.000  8.318987 6.452800
 04/04/2006 05:02:30.000        NA       NA
 04/04/2006 05:07:30.000  8.275098 6.493786

Here is Tsoil and Qsoil data from data file isff_20060404_120000.dat 

              (GMT)     Tsoil      Qsoil
2006 04 04 12:00:03.188 8.322      6.453          0          0      -9999
2006 04 04 12:00:08.176 8.322      6.453          0          0      -9999
2006 04 04 12:00:13.176 8.322      6.453          0          0      -9999
2006 04 04 12:00:18.176 8.323      6.453          0          0      -9999
2006 04 04 12:00:23.177 8.322      6.453          0          0      -9999
2006 04 04 12:00:28.177 8.322      6.453          0          0      -9999
2006 04 04 12:00:33.177 8.322      6.453          0          0      -9999
2006 04 04 12:00:38.177 8.301      6.453          0          0      -9999
2006 04 04 12:00:43.187 8.322      6.453          0          0      -9999
2006 04 04 12:00:48.177 8.322      6.453          0          0      -9999
2006 04 04 12:00:53.178 8.302      6.453          0          0      -9999
2006 04 04 12:00:58.178 8.343      6.453          0          0      -9999
2006 04 04 12:01:03.178 8.281      6.453          0          0      -9999
2006 04 04 12:01:08.170 8.343      6.453          0          0      -9999
2006 04 04 12:01:13.178 8.281      6.453          0          0      -9999
2006 04 04 12:01:18.170 8.343      6.453          0          0      -9999
2006 04 04 12:01:23.179 8.322      6.453          0          0      -9999
2006 04 04 12:01:28.179 8.302      6.453          0          0      -9999
2006 04 04 12:01:33.169 8.281      6.453          0          0      -9999
2006 04 04 12:01:38.169 8.322      6.453          0          0      -9999
2006 04 04 12:01:43.171 8.323      6.453          0          0      -9999
2006 04 04 12:01:48.172 8.301      6.453          0          0      -9999
2006 04 04 12:01:53.172 8.322      6.453          0          0      -9999
2006 04 04 12:01:58.170 8.343      6.453          0          0      -9999
2006 04 04 12:02:03.182 8.302      6.453          0          0      -9999
2006 04 04 12:02:08.162 8.302      6.453          0          0      -9999
2006 04 04 12:02:13.173 8.322      6.453          0          0      -9999
2006 04 04 12:02:18.171 8.302      6.453          0          0      -9999
2006 04 04 12:02:23.171 8.281      6.453          0          0      -9999
2006 04 04 12:02:28.171 8.302      6.453          0          0      -9999
2006 04 04 12:02:33.172 8.281      6.453          0          0      -9999
2006 04 04 12:02:38.172 8.302      6.453          0          0      -9999
2006 04 04 12:02:43.182 8.302      6.453          0          0      -9999
2006 04 04 12:02:48.172 8.322      6.453          0          0      -9999
2006 04 04 12:02:53.172 8.343      6.453          0          0      -9999
2006 04 04 12:02:58.173 8.343      6.453          0          0      -9999
2006 04 04 12:03:03.173 8.322      6.453          0          0      -9999
2006 04 04 12:03:08.173 8.302      6.453          0          0      -9999
2006 04 04 12:03:13.173 8.301      6.453          0          0      -9999
2006 04 04 12:03:18.173 8.302      6.453          0          0      -9999
2006 04 04 12:03:23.184 8.281      6.453          0          0      -9999
2006 04 04 12:03:28.174 8.281      6.453          0          0      -9999
2006 04 04 12:03:33.174 8.281      6.453          0          0      -9999
2006 04 04 12:03:38.174 8.322      6.453          0          0      -9999
2006 04 04 12:03:43.174 8.281      6.453          0          0      -9999
2006 04 04 12:03:48.175 8.302      6.453          0          0      -9999
2006 04 04 12:03:53.175 8.301      6.453          0          0      -9999
2006 04 04 12:03:58.175 8.281      6.453          0          0      -9999
2006 04 04 12:04:03.185 8.302      6.453          0          0      -9999
2006 04 04 12:04:08.175 8.323      6.453          0          0      -9999
2006 04 04 12:04:13.176 8.302      6.453          0          0      -9999
2006 04 04 12:04:18.176 8.281      6.453          0          0      -9999
2006 04 04 12:04:23.176 8.281      6.453          0          0      -9999
2006 04 04 12:04:28.176 8.301      6.453          0          0      -9999
2006 04 04 12:04:33.176 8.323      6.453          0          0      -9999
2006 04 04 12:04:38.177 8.302      6.499          0          0      -9999
2006 04 04 12:04:43.177 8.281      6.453          0          0      -9999
2006 04 04 12:04:48.177 8.322      6.499          0          0      -9999
2006 04 04 12:04:53.177 8.281      6.499          0          0      -9999
2006 04 04 12:04:58.177 8.281      6.453          0          0      -9999

148: daily status, Site all, Tue 04-Apr-2006 09:54:14 PDT, Daily status
Daily status April 4, 2006

On site: Horst and Semmer

All sensors operating normally at this time.  10m sonic and kryptons at west
affected by rain last night.

147: wireless-network, Site central, Tue 04-Apr-2006 09:50:39 PDT, lost rf connection to central
Lost wireless connection to central around 0600 PST this morning.  Briefly lost
connection to west while Steve trying to fix this problem and dsm server and 
netcdf process stopped.  Fixed connection to west and restarted dsm server and
netcdf process.  Steve recycled power on the etherant at central to restart 
around 10 am PDT.

146: radiation, Site south, Tue 04-Apr-2006 08:51:55 PDT, loggers at West and South
I changed the transmit power of the ZigBee units at west and south. They are now down to 10 dBm verses 18dBm. 
145: wireless-network, Site central, Tue 04-Apr-2006 07:57:43 PDT, Link to central down
The link to central has gone down. This occurred at about 6:00 am this morning.
The link to the hotfilm is still good so I do not think it is the TPOP;
however, I will recycle power on TPOP to see if that fixes it.
Power recycle did not help. Time for visit to central.

144: radiation, Site south, Mon 03-Apr-2006 15:50:26 PDT, Intermittent logger data
There is once again intermittent logger data showing up in the netcdf files and
Splus plots.  However, I have scanned an entire raw data file for a time when 
this occurs and see not missing data.
143: wireless-network, Site south, Mon 03-Apr-2006 15:42:17 PDT, replaced MaxStream radio in logger at south
REplaced 316 MHz MaxStream radio in logger with 2.4 GHz Zigbee radio.

Note that the intermittent logger data at south and west have returned.  We 
suspect a software problem at the base.
142: LOG, Site central, Mon 03-Apr-2006 15:39:48 PDT, retensioned guy wires
Tightened the N and SE, highest guy wires by two turns.  Could not tighten
the SW guy wire because the turnbuckle is fully compressed.

After tensioning and with a strong south wind:

N = 290 lbs
SW = 340 
SE = 320
141: HotFilm, Site central, Mon 03-Apr-2006 10:56:05 PDT, hot film procedures
Hot Film:
	The hot film system consists of 4 major components: (1) power source, 
(2) sensors (hot film and sonics), (3) data collection (Prometheus and Ripple), 
and (4) the communication system. The system can run without the communication 
link going however you will not be able to monitor on a real time basis.

     This assumes that power is off on all systems and all sensor cables are 
plugged in.

1: Power system
     The power system is designed to be running at all times. The power system 
consists of 2 parts; the batteries and the fuel cell. The fuel cell keeps the 
batteries charged.  Verify that the fuel cell system is running. If it is not 
operating then check the hydrogen tank. The tank pressure should be at least 
300 psi. If it is below 300 psi then it should be changed.  If you do change 
the tank remember what the setting of the regulator is  between 10 and 20 psi 
on the second stage. After the tank is attached then you can start the fuel 
cell. There is an on/off switch on the fuel cell located below the hydrogen 
line. Turn the switch on and the fuel cell should start. It will make a noise 
for about 15 seconds then start running.  If it stops you may need to replace 
the 9 volt battery located in the fuel cell. This requires removing the fuel 
cell from the enclosure. The battery is located on the side of the cell.

2:  Startup Ripple
     Ripple is located inside the electronics box attached to the top of the 
lid. Open the box and plug the power cable into one of the power ports on the 
12 volt battery splitter cable. Attach the USB drive if it is not connected. 
This is necessary in order for the aster processes to get going. Undo the 
bungee cord so you can open the DELL. BE CAREFUL UNHOOKING THE BUNGEE CORD. 
Power up the DELL and login when requested. 
Do a check_aster command to verify that the aster process is running. 
Also do a ping aster to verify that the wireless communication link is up.

3: motor controllers
     The motor control systems should be plugged into serial ports 9 (1m), 
10 (2m), and 11 (3m). With the lid open verify that the internal power cord to 
the Prometheus stack is unplugged. Now turn the main power switch on. This 
should start the 3 motor systems. The zero position sequence starts with the 
motor shaft moving CW to about 135 degrees. It stops for a second then goes 
CCW to the zero position. Zero is parallel to the sonic boom direction. You 
want to make sure all three probe support arms follow the sequence above. If 
something does not go right then turn off the main power switch immediately.
NOTE: At this time NO probes should be mounted.

4: Tri-axial probes
	The tri-axial probes can be installed. Note the reference number on 
each probe box. The probe order is: #1 – 1m, #2 – 2m, #4 – 3m. When 
inserting the probes note the notch cut out at the back end of the probe. 
This should be aligned with a corresponding piece on the probe support. 
Be extra careful handling the probes!  Put the yellow protection cap, 
on the end of the support arms into the probe box.
5: Prometheus
	The Prometheus can now be plugged in. You can run minicom on ripple 
if you want to watch it boot up. After a minute or two you should see a cockpit 
come up on ripple. Verify everything is running as expected. Also verify data 
is being written to the USB drive. This can be done by either watching for a 
green light flash or do a DF command on ripple and verify that it is increasing.
At this time the hot film data on the analog channels should be around zero 

6: CTA power
	Now plug in the CTA power cable. It is plugged into on of the aux ports
on electronics box.  Move it over to the other aux port. You should be getting 
A/D on the cockpit.

	Shutdown is pretty much in the reverse order as startup

1: CTA power
	Unplug the CTA power cord and move it back to the aux port.

2: Prometheus
	Via minicom or ssh to daisy, do a halt command to stop the Prometheus.

3: Tri-axial probes
	Remove the hot film probes. Do them one at a time to ensure they get 
put back in their correct box. 
Remember to put back on the yellow protection cap.

4: Ripple
	Ripple can now be shutdown. Once it is shutdown remove the USB drive so
that the data can be downloaded back at the base. 
Remember to unplug the power cable and put it back inside the box.

Other items of interest:
Sonics: the sonics are routed to 2 locations: Prometheus and central. 
Only the transmit signal is routed to the Prometheus. If you ever have to 
disconnect the sonic cables from the Prometheus box make sure you note how 
they are connected.
Sonic 1m – port 5
Sonic 2m – port 8
Sonic 3m – port 7

Ripple: Ripple may be running. Tests have shown that the fuel cell is apply to 
keep ripple up all the time if its internal battery is fully charged. However, 
if the USB drive is not attached the aster processes may not be running!

140: daily status, Site all, Mon 03-Apr-2006 09:35:24 PDT, Daily status
Daily status April 3, 2006

On site: Horst and Semmer

All sensors operating normally at this time.

139: daily status, Site all, Sun 02-Apr-2006 11:45:16 PDT, Daily status
Daily status April 2, 2006

On site: Horst and Semmer

All sensors operating normally at this time.
Hot-film EOP last night.

The periodic losses of data from the loggers at south and west no longer appear
in the netcdf data, even for prior time periods!?!?!?

The wireless network went down yesterday, but data continue to be recorded 
at the tower sites.

138: base, Site base, Sun 02-Apr-2006 11:08:32 PDT, how to restart covars
To restart covar/netcdf process:

[aster@aster ~]$ cd ~aster
[aster@aster ~]$ ./runstats_sock
18612: old priority 0, new priority 19
[aster@aster ~]$ ps -ef |grep stats
aster    18615     1  3 10:49 ?        00:00:00 statsproc sock:localhost:30000

To catch up covar/netcdf in ~aster edit for days to rerun, e.g.

statsproc isff_20060402_*.dat


[aster@aster ~]$ batch
at> ./
at> D
job 18 at 2006-04-02 10:52
[aster@aster ~]$ at -l
18      2006-04-02 10:52 = aster
[aster@aster ~]$ ps -ef | grep stats
aster    18615     1  2 10:49 ?        00:00:07 statsproc sock:localhost:30000
aster    18642 18641  0 10:52 ?        00:00:00 /bin/sh ./
aster    18644 18642 93 10:52 ?        00:02:23 statsproc isff_20060401_000000.dat isff_20060401_040000.dat isff_20060401_080000.dat isff_20060401_120000.dat isff_20060401_160000.dat isff_20060401_221636.dat isff_20060401_224518.dat

137: krypton, Site west, Sun 02-Apr-2006 08:12:47 PDT, Krypton clean and s/n at West
The kryptons were cleaned at west at ~16:45 pm.

kh2o s/n at 5 m = 1395
kh2o s/n at 30 m = 1258
136: HotFilm, Site central, Sun 02-Apr-2006 08:04:11 PDT, Hot film status
The hot film system was up and running around 16:00 yesterday. The fuel cell
system was able to keep ripple going since it was powered up for Gordon on
The deadband was set to 20 degrees.

135: radiation, Site , Sun 02-Apr-2006 08:00:17 PDT, Rad system ar west running
Tom notice that the 2.4 Ghz radios did not solve the missing radiation data.
So after getting the hot film going, we took the extension serial cables for
the hot film sonics to hardwaire in the logger at west. This took place around
16:30 yesterday.
134: sonic, Site central, Sat 01-Apr-2006 15:11:28 PST, 25 m sonic at central
Restarted 25 m sonic with rserial -h /dev/ttyS10 central.
Then & T Ae D & (?)
133: wireless-network, Site base, Sat 01-Apr-2006 14:45:57 PST, restared wireless network and dsm server
Spent the morning trying to restart communications with central with no luck.
Then also lost communications with west and south and the dsm server stopped
(do not know in what order this occurred).

Could not restart dsm server until communications were restored.  Steve
recycled power on tpop at base, restored communications to all three sites,
and then the dsm server was able to be restarted.

Wireless link to sodar is intermittent.

132: krypton, Site central, Sat 01-Apr-2006 13:48:14 PST, maintenance at central
Steve and Tom at central tower.  Steve climbed 30 m tower 11:40-12:20 PST to
investigate communication problem and clean kh2o at 30 m.  Then cleaned kh2o
and Licor 7500 at 5 m; off 5 m tower at 12:34 PST.

KH2O s/n = 1394 @ 30 m
           1390 @ 5 m

Finally, Steve cleaned 10 m radiometers with mop.
Also added yokes for steps and safety rope on 10 m radiometer tower.
131: radiation, Site west, Fri 31-Mar-2006 12:26:46 PST, logger at west
The Campbell data loggers at south and west have been losing data intermittently
which Steve suspects may be due to 915 MHz interference.
He removed the logger at west around 11:45 PST to replace the MaxStream radio 
with a Zigbee radio.

   Went to re-install logger at west at about 16:00. For some unknown reason the cr10 lost its program. Will fix tomorrow.

Update 4/1/06:  Steve went to reprogram logger and after switching cables to do this, it started on its own, so apparently did not lose its program as thought.

130: daily status, Site all, Fri 31-Mar-2006 10:12:57 PST, Daily status
Daily status March 31, 2006

On site: Horst and Semmer

All sensors operating normally at this time.

The Campbell data loggers at south and west have periodic data losses.
Steve suspects that this may be interference at 915 MHz.

129: krypton, Site all, Thu 30-Mar-2006 12:55:06 PST, Krypton cals & location
According to logbook entry #9, here's the krypton layout:

     5m     1390
    30m SN: 1394

     5m SN: 1391
    30m SN: 1393

     5m SN: 1395 (no cal below)
    30m SN: 1258


1/19/2006	K1258
Calibration	TREX06
pathlength	1.3
Kw		-0.1576957
V0		4306.955643

1/21/2006	K1390
Calibration	TREX06
pathlength	1.3
Kw		-0.1513344
V0		968.0184596

1/21/2006	K1391
Calibration	TREX06
pathlength	1.3
Kw		-0.1511835
V0		3961.417306

1/19/2006	K1393
Calibration	TREX06
pathlength	1.3
Kw		-0.1584523
V0		2988.347174

1/21/2006	K1394
Calibration	TREX06
pathlength	1.3
Kw		-0.1656531
V0		2601.645432

/19/06		K1395
Calibration	TREX06
pathlength	1.3
Kw		-.1551411
V0		4295.005517

1/21/2006	K1397
Calibration	TREX06
pathlength	1.3
Kw		-0.1275433
V0		1723.074647

1/19/2006	K1525
Calibration	TREX06
pathlength	1.3
Kw		-0.162492
V0		6556.380341


128: soil, Site all, Thu 30-Mar-2006 12:27:01 PST, gravimetric soil moisture
gravimetric soil moisture measurements

South at 11:45 PST

3-6 cm	total mass = 176.4 - 73.1 gm = 103.3 gm
	wet (gm)	dry
d1	19.9		19.2
d2	27.9		26.9
d3	21.3		20.5

0-3 cm total mass = 154.0 - 45 gm = 109 gm

e1	20.4		20.1
e2	20.7		20.3
f2	19.0		18.7

Central at 14:40 PST
Soil samples were too dry to provide a valid bulk density

3-6 cm

b2	17.8		14.5
b3	16.0		13.0

1-3 cm

c1	11.0		9.8
c2	11.7		10.4

West at 15:40 PST
Soil samples were too dry to provide a valid bulk density

3-6 cm

a1	19.6		18.7
a2	18.3		17.4

1-3 cm

a3	26.8		26.2
b1	24.7		24.0

127: HotFilm, Site central, Thu 30-Mar-2006 11:37:01 PST, HotFilm shutoff
The hotfilm was shutdown at about 10:45am local. The sonics were moved over
to central nids.
The 35 watt fuel cell kept things going and the batteries still are up
around 12;5 volts.

126: soil, Site south, Thu 30-Mar-2006 11:04:49 PST, Qsoil negative at south
Qsoil reads about -2.5% at south.
125: daily status, Site all, Thu 30-Mar-2006 10:47:15 PST, Daily status
Daily status March 30, 2006

On site: Horst and Semmer

All sensors operating normally at this time.
15 m sonic had to be rebooted last night.
Hot-film EOP started last night at 16:35 PST; ran okay through the night.

124: HotFilm, Site central, Wed 29-Mar-2006 16:55:35 PST, Hot Film status
The hot film was offically back on the air at 16:35.
During the process the 1m, 2m, and 3m sonics were removed from nids central
and move back to daisy. An option to eliminate this cable changing is to
build special Y cables to route the sonic xmt line into daisy. Maybe in the

123: radiation, Site central, Wed 29-Mar-2006 16:53:13 PST, Logger #1 at central
The logger # 1 at central is back up (~16:15). The code changes to turn on
and off the ZigBee did not work. The transceiver comes on but the receive
light does not. The logger was put back into tis old mode.

122: LOG, Site all, Wed 29-Mar-2006 11:44:56 PST, guy wire tensions
Measured guy wire tensions today, March 30:

	se	ne	w
1	395	350	360
2	330	285	300
3	395	300	290
4	350	300	300

1	430	460	435
2	330	320	350
3	300	345	315
4	305	295	310

1	400	355	420
2	350	300	300
3	400	280	350
4	340	300	310
121: radiation, Site central, Wed 29-Mar-2006 10:18:24 PST, Shutdown pyg sensor at central, logger ID 2
I have removed logger ID 2 from central to make some modes to the software
so that the radio will go to sleep between data transmissions. This should
help the power issue at the two single pygreometer sites, The logger was
removed around 9:15 am local,

120: daily status, Site all, Wed 29-Mar-2006 09:40:08 PST, Daily status
Daily status March 29, 2006

On site: Horst and Semmer

All sensors operating normally. 
Yesterday we added the 3 hot-film sonics to the normal data collection.

119: HotFilm, Site central, Tue 28-Mar-2006 14:58:41 PST, Hot film status
The hot film system was left up overnight to collect sonics data.
Unfortunately the batteries could not keep things up and the system went down.
Earlier this morning I changed the batteries and brought the system back up.
I re-routed power for daisy to come from the nids power source. The fuel cell
system will now provide power for ripple only.

The sytem was taking down once we got the sonics interfaced to nids (refer to
messages 117 and 118). We will see if the fuel cell keeps the batteries
charged at hot film.

118: sonic, Site central, Tue 28-Mar-2006 14:55:57 PST, Hot film sonics on central nids
The 3 hotfilm sonics are now going through nids central. This was done
by extending the sonic cables to each the nids box. Whenever the hot film
is active these sonics need to be reconnected to Daisy.
117: sonic, Site central, Tue 28-Mar-2006 13:19:05 PST, hot-film sonics added to NIDS
In order to collect the 1,2,3 m sonic data when the hot-films are not running,
we are cabling them to the viper data system.

I have saved ads3.xml as ds3.xml.old and modified ads3.xml with the
following additional code:


116: daily status, Site all, Tue 28-Mar-2006 10:00:01 PST, Daily status
Daily status March 28, 2006

On site: Horst and Semmer

All sensors operating normally. 

115: HotFilm, Site central, Mon 27-Mar-2006 15:42:17 PST, HF site back up collecting sonic data at 15:30
The hotfilm site was brought back on the air collecting sonci data only.
This was done vias channel_config. There are now wo channel_configs in
OPS. "channel_config.sonci" is used for only sonic data while
"channel_config.hf" is the hotfilms added. You have to copy either of
these files to "channel_config" before starting things up.

The battery voltage levels were down around 11.75 volts. The hydrogen
tank was at ~1800 psi. This will be a good test to see if the fuel cell
can recharge the batteries wihtout the hot film stuff going. I also
disconnected power to the CTA moduels.

114: data-system, Site west, Mon 27-Mar-2006 14:13:29 PST, Copy local storage to aster over the network
Created a new directory on /mnt/exthd8/projects/TREX called local.
Downloaded data with, e.g.
scp root@west:/var/tmp/usb/TREX/west_20060326_000000.dat 
113: radiation, Site central, Mon 27-Mar-2006 11:42:15 PST, Battery changed at logger #1 at central
The battery was changed at radiaiotn logger #1 at central.
This took place around 10:45 this morning.

112: HotFilm, Site central, Mon 27-Mar-2006 11:26:34 PST, Hot film status
The hot film ran through out the night until ~10:30 this morning. The data
are bieng downloaded to exthd7 in the base. Once the pocketec drive is
cleared off it will be takne back out to central where the system will be
turned on to collect sonic data only.
NOTE: the battery level was down to 12 volts. The fuel cell was running
fine so this down time will be good for the batteries.

111: daily status, Site all, Mon 27-Mar-2006 10:05:24 PST, Daily status
Daily status March 27, 2006

On site: Horst and Semmer

All sensors operating normally.  Hot film system running since 1500 yesterday.

110: HotFilm, Site central, Sun 26-Mar-2006 14:10:07 PST, Deploying hot film probes
Deployed all three hot film probes this afternoon for a night run.

Data started about 1500 pst.

109: daily status, Site all, Sun 26-Mar-2006 10:44:16 PST, Daily status
Daily status March 26, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the hot film system.
Visited west yesterday to install new battery for electric fence.

dsm server went down yesterday evening and PAM mast with etherant antenna to
central broke in apparent high wind.  Don't know if these events were related.

hot film motor orientation appears to have run successfully last night for
all three units.

108: data-system, Site base, Sun 26-Mar-2006 09:47:08 PST, dsm server down
dsm server stopped yesterday at 0225 GMT, March 25; restarted this morning.

107: wireless-network, Site base, Sun 26-Mar-2006 09:33:24 PST, PAM mast broke last night
The PAM mast broke in an apparent high wind event last night; guy wire 
turnbuckles were not tied-off with cable ties!

Lost communication with central.

Repaired mast this morning at ~0930 PST.  Steve restarted etherant communicationby recycling power at central tower.

106: HotFilm, Site central, Sat 25-Mar-2006 17:04:09 PST, hotfilm update
Mods have been completed on the motor controller circuit. All 3 units are
now deployed at central. The plan is to do a run tonight with breakable
wooden sticks in place of the probes. If the motor controllers go beyond
the limit switches the wooden probes should break.
If things run goo over night then we are back in business to collect hotfilm

105: LOG, Site west, Sat 25-Mar-2006 16:38:08 PST, replaced fence battery at west tower
Since we saw cows yesterday near the west tower, I put a fresh battery on the
104: daily status, Site all, Sat 25-Mar-2006 09:16:13 PST, Daily status
Daily status March 25, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the hot film system.
Visited central site yesterday to restart etherant communications;  
visited west tower to clean KH2O.

Steve continues to rebuild the motor orientation mounts for the hot-film to
correct slipping problem.

103: HotFilm, Site central, Fri 24-Mar-2006 17:30:25 PST, hotfilm status
Work was spent on the motor controller systems. SInce one of the optic units
was not acting correctly I decided to re-wire it to a different format. There
is now a pull down resistor on the emitter side of the detector. The
result was very positive. The signal levels pull all the way to 0 volts as
wanted. I will need to re-write the code because the trigger logic is now
reversed; ie, it triggers on a falling edge verses a rising edge. I should
get this done tonight so that we can deploy them tomorrow.

102: krypton, Site west, Fri 24-Mar-2006 14:41:30 PST, Cleaned KH2O at west
Steve cleaned the KH2O at west; on tower from 1258 - 1318.

Cows in vicinity.  Need to activate the fence.

101: wireless-network, Site central, Fri 24-Mar-2006 10:51:16 PST, data link to central fixed
The data link to central went down around mid-day yesterday; when Steve was
climbing tower??

Went to central this morning & recyled power on both the etherant and the viper.
Link working again.
100: daily status, Site all, Fri 24-Mar-2006 10:41:27 PST, Daily status
Daily status March 24, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the hot film system.  
Visited west tower yesterday for routine maintenance; replaced fuse on 
serial channel 14 and restarted aerosol spectrometer.

Steve has been rebuilding the motor orientation mounts for the hot-film to
correct slipping problem.
99: HotFilm, Site central, Fri 24-Mar-2006 07:51:04 PST, Hotfilm work from yesterday
The hotfilm system was shut down yesterday. I brought the motor controller
systems back to the base. We are having a problem with the coupler between
the motor shaft and the vertical probe support shaft. Also problems with the
coupler and the optics units.

98: radiation, Site west, Thu 23-Mar-2006 15:43:15 PST, maintenance at west site
March 23: Tom at west site for maintenance; arrived ~ 1415 pst

Guywire tensions (lbs)
1	450	360	375
2	300	290	310
3	385	300	280
4	340	295	290

Cleaned radiometers with methanol 14:38
Replace fuse on serial channel 14 to restart aerosol spectrometer.
Alone, so did not clean KH2O
97: radiation, Site central, Thu 23-Mar-2006 13:47:19 PST, maintenance at central site
March 23: Steve and Tom
Service visit to central site
Guywire tensions:
1	380	450	420
2	325	310	350
3	310	370	300
4	320	300	305

Steve cleaned KH2O and LI-7500 with methanol
Steve on 30 m tower 11:35 - 12:05
Steve of 5 m tower  12:05 - 12:10

Then cleaned radiometers with methanol spray, wiped off radiometers
with kleenex.

96: radiation, Site south, Thu 23-Mar-2006 10:38:56 PST, Maintenance at south yesterday
Steve and Tom visited south tower Mar 22 before noon for routine maintenance.
Measured guywire tensions, all above 300 lbs:

1	400	370	410
2	360	325	320
3	400	320	340
4	340	330	315

Steve cleaned radiometer domes and KH2O windows with methanol.
He wiped off the net radiometer domes.
All radiometers appeared fairly clean prior to cleaning.

We crashed NIDS by running rserial locally.  This is a known
bug.  Should have run through ASTER (ssh aster@server)

95: daily status, Site all, Thu 23-Mar-2006 09:55:58 PST, Daily status
Daily status March 23, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the aerosol sensor
at west and the hot film system.  Did not get to west yesterday to recycle 
power; maybe today.

Visited south tower yesterday for routine maintenance.

Mounted hot film sensors yesterday around 1700 yesterday, 
pointed in a fixed direction (west) because motor orientation device slipping.
Ran until about 3:45 this morning when daisy appears to have stopped.

94: HotFilm, Site central, Thu 23-Mar-2006 07:45:27 PST, hotfilm
Based on the archive files on ripple, diasy shut down around 3:45 this morning.
I assume that the reset program "ndaqrestart" used in the crontab may have a
bug. I will talk to Gordon.

I plan to go out there in teh next 1/2 hour to shut the system down and bring
back the USB drive.

93: base, Site base, Wed 22-Mar-2006 20:36:58 PST, Network disconnected
Charlie called Cebridge and found out that they had disabled our modem,
After a few minutes they activated it again. It may get disabled again
so we will leave it on the school link until tomorrow.

92: base, Site base, Wed 22-Mar-2006 19:35:16 PST, CRASH!
Around 7:00pm we had a major crash. The first thing to go down was the
network interface. Charlie believes it was a Cebridge problem. I then
notice that the aster system was not working correctly. AA hard reboot of
aster had to be done. It refused to allow su login. It would not come
up properly, hung at swap disk space.Recycled power again and it came up ok.
Charlie is still working on the network link.

91: HotFilm, Site central, Wed 22-Mar-2006 17:44:30 PST, htofilm - probe position problem
All shafts on probe position system have loosened up. I believe the problem
is the tolerences between the optics and the shaft coupler along with the pin
used to trigger the optic. I have seen two problems:
(1) The shaft coupler piece has had contact with an optic detector causing
the detector to move out of position.
(2) The roll pin has made contact with the optic detector as it passes through.
The contact point has been at the bottom of the optic detector.

Possible solution:
(1) reduce the diameter of the shaft coupler towiden the space between the
coupler and the optics.
(2) change the roll pin to a flat surface piece. This would provide more
clearence between the top and bottom of the optic piece.
90: HotFilm, Site central, Wed 22-Mar-2006 17:31:54 PST, Installed hot film sensors
Installed hot film probes and began running at ~1712 pst.

Had problems with orientation shafts slipping, so turned off orientation motors
and oriented probes ~ west.

1 m probe had problem with y bnc connector immediately after probe.  Steve
jiggered with it and got it connected for now. I believ this is the connector
that John had found to be bad.

The batteries were changed. The old olds were done to ~12 volts.The fuel
cell should be able to recharge the batteries during the daytime when
the system is off. 

Probe 1 at 1 m
Probe 2 at 2 m
Probe 4 at 3 m

89: krypton, Site all, Wed 22-Mar-2006 10:05:43 PST, Cleaning KH2O windows
I called Campbell today and they recommended muriatic acid for cleaning KH2O

88: daily status, Site all, Wed 22-Mar-2006 09:41:46 PST, Daily status
Daily status March 22, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the aerosol sensor
at west and the hot film system.  Will try recycling power on aerosol sensor.

Hot film system ran successfully last night without hot film probes attached.
Approximately 0.5 hours of A/D data was lost due to failure of daisy A/D,
but daisy was rebooted once per hour to restart A/d. 

87: HotFilm, Site central, Tue 21-Mar-2006 16:55:56 PST, hotfilm
   I believe the problem with ripple has been found. It may have been in the USB
port connection. The PCMCIA USB card is a USB2 connection while the USB
connection on the back of ripple is a USB1. I changed to the ripple USB at
about 11:30 this morning and it has been running fine every since.

  Prometheus is still having problems. It only wants to stay up for a few
minutes. The plan now is to install a forced reboot on Prometheus every xx
minutes where xx will be determined based on how it is running.

   The fuelcell is working like it was meant to be out there. It has used about
100 PSI over the last 24 hours.

86: HotFilm, Site central, Tue 21-Mar-2006 10:11:27 PST, hotfilm status
During the night ripple shut down so no data was collected. The 35 watt
fuel cell kept runnin gok and the batteries are charged. I went out this
morning to restart ripple. I also remounted the 1m motor system. I had to
remove the 2m motor system becasue the motor shaft coupler screws were loose.

85: daily status, Site all, Tue 21-Mar-2006 09:40:50 PST, Daily status
Daily status March 21, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the hot-film system.

Still struggling with hot film system.  Test last night failed.

84: LOG, Site base, Mon 20-Mar-2006 16:51:26 PST, How to restart the dsm_server
If you get a response of connection refused to data_stats sock:localhost,
check ps -ef |grep dsm_server. If not running, log in as root and run
/etc/init.d/nids restart. This should get the dsm_server running again.

NOTE: if the wireless link is down, dsm_server will not get going.

83: HotFilm, Site central, Mon 20-Mar-2006 16:47:31 PST, Htofilm status
During the day the hotfilm was left running at central. The A/D did not stay
up all the time. Back at the base I got the 35 watt fuel cell running. It 
was taken out to central to serve as the main power source for the hotfim
stuff. Tis included ripple which is mounted inside the Prometheus box.
In addition Gordon made some changes so that the USB harddrive is now
interfaced to ripple instead of the Prometheus. The system was back on the
air at about 16:40 local time.

82: wireless-network, Site , Mon 20-Mar-2006 10:22:30 PST, Recycled power at school
Network to school down this morning.  Recyled power in attic and back up.
81: krypton, Site all, Mon 20-Mar-2006 10:21:24 PST, KH2O serializer data
FYI, the "serializers" don't digitize fast enough to keep up with 60Hz, 
so every fifth(?) sample from the kryptons is NA.  Since humidity fluxes 
are a low priority for TREX, and fluxes can still be calculated, I have 
not been worried about this .

...Steve Oncley

80: daily status, Site all, Mon 20-Mar-2006 09:21:05 PST, Daily status
Daily status March 20, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the hot-film system.
The link to the school is down at the moment.

The hot film system stayed up during the night. Based on the local storage
file sizes the A/D stayed up for about 7 hours. Ripple stayed up on an
external battery that is not on the charging grid. We will go out there
later this morning to check on the ripple battery and restart the A/D system.

79: HotFilm, Site central, Mon 20-Mar-2006 07:30:27 PST, hotfilm
The system stayed up during the night. Based on the local storage file sizes
the A/D stayed up for about 7 hours. Ripple stayed up on an external battery
that is not on the charging grid. I will go out there later this morning to
check on the ripple battery and restart the A/D system.

78: power, Site central, Sun 19-Mar-2006 18:35:04 PST, fuel cell
The 60 watt fuel cell was at 14.68 volts and 800 psi at 6:15pm.

77: HotFilm, Site central, Sun 19-Mar-2006 18:29:42 PST, hotfilm status
The hotfilm was moved out to central this afternoon. There were some initial
crashes on the A/D. After some playing around it seemed to stay up. The 3m
sonic was not responding correctly. Who knows if this may have been having
some effect. It was disconnected. Ripple is running off an external battery
via the car charging unit that Steve Oncley had. There are no hotfilm probes
deployed. Also the motor controller at the 1m height was acting up. It was
not seeing the zero position switch. It appears to be a similar problem we
were having back in Boulder. What is real frustrating is that unit was in the
base during the tests over the last few days and it worked without any

76: daily status, Site all, Sun 19-Mar-2006 13:18:28 PST, Daily status
Daily status March 19, 2006

On site: Horst and Semmer

All sensors operating normally, with the exception of the hot-film system.

The hotfilm stayed up last nightin the lab. It was configured in the base 
with ripple tied directly to the prometheus via a swapper ethernet cable. 
The plan is to move it out to the site later today with ripple housed 
inside the Prometheus box. The external hard drive on ripple will not be used 
since the local USB drive files look good. We may try to get ripple hooked up 
to the wireless link so we can monitor things back at the base.

75: HotFilm, Site base, Sun 19-Mar-2006 08:44:25 PST, HotFilm status
The hotfilm stayed up last night. It was configured in the base with ripple
tied directly to the prometheus via a swapper ethernet cable. The plan is
to move it out to the site later today with ripple housed inside the
Prometheus box. The external hard drive on ripple will not be used since
the local USB drive files look good. We may try to get ripple hooked up to the
wireless link so we can monitor things back at the base.

74: power, Site central, Sat 18-Mar-2006 14:15:48 PST, central fuel cell
At noon the fuel cell was reading 15.3 volts and the pressure was at 1050.

73: LOG, Site central, Sat 18-Mar-2006 14:13:47 PST, Central visit
Tom and I went to central to do a few tasks. This included getting a soil sample, relocating the solar panel a logger 1 and checking the fuel cell. This took place around noon.

72: daily status, Site all, Sat 18-Mar-2006 08:40:51 PST, Daily status
Daily status March 18, 2006

On site: Horst and Semmer

Replaced dead battery in both loggers supporting the downward-looking 

Brought hot film to base for testing; ran for 4 hours last night until the
a/d failed.

All sensors operating normally, with the exception of the hot-film system.
71: power, Site central, Sat 18-Mar-2006 08:37:34 PST, Fuel Cell at central
I did a check on the fuel cell this morning.
At 8:20 voltage = 14.80 gas = 1100

70: HotFilm, Site base, Sat 18-Mar-2006 08:00:32 PST, hot film run in base
The hot film ran for about 4 hours in the base then the A/D died. I will restart daisy for another run.

69: HotFilm, Site central, Fri 17-Mar-2006 17:49:31 PST, HotFilm
Due to the discovery of the daisychain on the sonics, any data looked at since this time should be done with great care. In fact any sonic data representing 2m and 3m should be ignored.
68: HotFilm, Site central, Fri 17-Mar-2006 15:15:53 PST, Removed hot film data system and 1 m sonic
Steve S and Tom removed the hot film data system and 1 m sonic from central
site this afternoon to work on it in lab.

Steve discovered that the 2 and 3 m sonics were never cabled within the data 
system.  The 3-sonic replicator cable was still transmitting the 1 m sonic 
data to all three channels.It was also discovered that serial channel 201 was
not responding to a sonic input. So a change was made in channels.txt on daisy
(/home/isff). Channel 201 was moved from /dev/ttyS5 to /dev/ttyS7

Later that evening:
  Steve set the Hotfilm system up in the trailer. It has been running for
~1hour with one sonic and 9 analog at 2000 hz. It is now runing with 3
sonics and 9 analog. It will be left up over night to see what happens.

67: power, Site central, Fri 17-Mar-2006 10:59:42 PST, 60 watt fuel cell
I checked the output of the 60 watt fuel cell around 10:k00 am. It was at about 13 volts. So just for fun I tiurned down the fan power and the cell jumped up to about 15.5 volts. I plan on going out there at sundown to monitor it.

66: radiation, Site central, Fri 17-Mar-2006 10:55:58 PST, single pyg units
Spent some time out at central changing the batteries and re-orienting the solar panels at the two single pyg sites. This took place between 9:00 and 10:30 am.

65: daily status, Site all, Fri 17-Mar-2006 08:18:52 PST, Daily status
ISFF daily status, March 17 am

Staff:  Steve Semmer arrived yesterday, Steve Oncley left this morning
        Tom still here....

ASTER data directory filled up last night, so lost some data

One of the 4 data loggers at Central has stopped; likely dead battery.
This is one of the loggers supporting a single outgoing pyrgeometer.

Installed aerosol spectrometer at west tower for REAL yesterday.
Replaced bad 15 m TRH at south tower.

Took soil samples for gravimeteric soil moisture analysis at south and 
west towers.

64: LOG, Site west, Thu 16-Mar-2006 15:00:53 PST, west visit
- mounted MetOne aerosal counter at 15m (TRH is at 14.5m).  Returned at 14:30
to complete cabling
- took gravimetric soil sample.

63: LOG, Site south, Thu 16-Mar-2006 14:58:19 PST, South visit
- Replaced faulty 15m TRH (704) with (tested and working) spare 008.  This one
is working.
- Recleaned krypton which again lowered voltage from 0x5a00 to 0x4f00!  We'll
need to get new, fresh water and swabs and try again.
- Took soil sample (see comment #60)

62: LOG, Site central, Thu 16-Mar-2006 14:56:48 PST, Central visit
- Replaced empty H2 cylinder with full.
- Cylinder we're running on is at 1360psi -- about twice the usage compared to
the last 2 nights.

61: LOG, Site all, Thu 16-Mar-2006 14:36:29 PST, one-time tasks
Just a brief task list, to aid the transition from me to Semmer:

- can remove parabolic antenna, but left in place in case the signal to south
deteriorates (now 18db, was 40db with parabolic) and we don't need the 180deg
antenna for the school connection.
- unused 30m etherant to bulgin cable is still strapped all the way up the 
- partially-used hydrogen bottle can be removed (unless power at west goes
bad, now that we've added Shane's extra 6W -- but we just decreased station
power by ~3W by installing the beacon photocell)
- electric fence is almost completely dead.  If you want it running, need to
replace battery (at least).  It isn't clear that it is needed anyway, though.

- maxstream 915M radios should be replaced by 2.4G radios (but need to be
configured first).

- should add code to CR10x to power down radios when not transmitting.  Central
a1&a2 batteries are not doing well and will need to be replaced.
- need to shoot boom angles when theodolite is repaired

- chem shelter could be removed
- 15m sonic doesn't recover on power up.  Needs "&", "T", "??", "D" typed at
it through rserial.  [we discussed checking settings, but haven't yet done]
- one H2 cylinder (being used) is about half full, other is completely full.

- 3 empty sonic boxes could be moved to REAL seatainer
- big winch under steps could be moved to REAL seatainer (for security)
- "Data tapes" plastic case on shelf has 2 TRHs to be repaired.  One is 
probably a bad 50Y connection.  The other probably is a dead fan.
- have spare csat3 from Inyo County (via Greg), but not wired for us (power
through 232 connector).
- printing doesn't work through aster docking station.  Have connected printer
to ripple, but won't work if ripple is at central.

60: soil, Site all, Thu 16-Mar-2006 14:27:09 PST, gravimetric soil measurements
Gravimetric samples taken at south (~1130) and west (~1345) today.  Wet 
weighing done ~1420 -- a bit late for south, but lets hope...

3/18/06: Samples taken at central tower (~1215 pst), weighed at ~2030 pst

Divided into 2 layers: 0-3cm and 3-6cm, 3-6 being roughly where the Ech2o
probes are installed.

Numerous small rocks in west sample.

Total weight (gm):

Central: 0-3cm: 89.4 - 45.0
         3-6cm: 175.2 - 79.2
South: 0-3cm: 126.8-26.9
       3-6cm: 166.0-53.7
West:  0-3cm: 141.5-26.8
       3-6cm: 141.9-48.4

Aliquots:      		wet

Central 0-3cm: d1-3	15.8	16.6	9.2
        3-6cm: c1-3	28.7	28.4	21.1
South:	0-3cm:  a1-3	38.2	34.3	27.2
     	3-6cm:  b1-3	39.2	37.0	35.0
West:	0-3cm:  c1-3	48.9	37.8	27.9
      	3-6cm:  d1-3	22.1	36.8	33.5	

Aliqouts:		dry

Central 0-3cm  d1-3	13.2	13.9	7.8
        3-6cm  c1-3	22.7	22.5	16.8

South:	0-3cm:  a1-3	37.3	33.6	26.4
	3-6cm:	b1-3	37.5	35.4	33.6
West:	0-3cm:	c1-3	48.3	37.2	27.3
	3-6cm:	d1-3	21.1	35.2	31.9

59: HotFilm, Site central, Thu 16-Mar-2006 09:31:00 PST, prometheus config change
As yet another test, I've just changed the config to enable sampling of only
1 triple hotfilm probe at 1000 and all 3 sonics.  The usb disk didn't come up.
I had tried to set the sample rate to 500 on all hotfilms, but the prometheus
went into a continuous ndaq restarting cycle -- perhaps 500 sps is not a valid

I had to start archiving manually, which is now being done to ripple's disk.
I figured that we might as well be acquiring data from 3 sonics, since they
don't significantly add to the data rate.

P.S. Even this test was a failure!  It ran for only 20 minutes and died.  
Data rates in the file were ok (no dropouts -- frequency as expected).

58: daily status, Site all, Thu 16-Mar-2006 09:04:02 PST, Daily status
Daily status March 16, 2006

On site: Horst and Oncley

Replaced 15 m hygrothermometer at south tower yesterday, but it is also
not operating properly.  One guess is that the humitter is not plugged in.

Soil moisture sensor at south tower reporting negative soil moisture.

Otherwise things appear to be normal for the moment.

57: power, Site central, Wed 15-Mar-2006 14:45:57 PST, Service visit to Central tower
Tom and Steve visited the Central tower after the West tower.

Check guywire tensions:

SE upper guys: upper guy = 300 lbs	325 lbs
               lower guy = 320        	315 
N  upper guys: upper guy = 280		300
               lower guy = 370		370
SW upper guys: upper guy = 300		325
               lower guy = 325		350

Second reading above is after increasing the tension on the uppermost guy wire
by one full turn.

H2 pressure = 2000 psi

Steve on tower cleaning 30 m KH2O: 12:05 - 12:20 PST
Steve on 5 m tower cleaning 5 m KH2O: 12:25 - 12:30 PST

Apparently cleaning the 30 m KH2O made little change in the output, but cleaning
the 5 m KH2O decreased the output from 0.9 V to 0.7 V !!

56: soil, Site west, Wed 15-Mar-2006 14:39:23 PST, Service visit to West tower
Tom and Steve visited West tower after South tower.

Replaced TP01 power cable within logger.  Data look good at trailer.
Steve cleaned KH2O at 5 m and 30 m.  On tower from 11:05 PST until
11:25 PST.

Apparently, cleaning the 30 m KH2O decreased the output from 1.7 V to 0.8 V and decreased the 5 m output from 2.9 V to 1.5 V !!!

55: TRH, Site south, Wed 15-Mar-2006 14:33:46 PST, Service visit to South tower
March 15, ~9.30 am - ~10:30 am

Steve and Tom visited South tower this morning.
Replaced blown fuse on NIDS serial channel 8 - 15m TRH.  Replaced fuse but Stevefound the fan was not working - probably blew the fuse again.  (Turned off 
station to replace fuse).  Replaced TRH with s/n 702(?) and replaced fuse
again - fan is working.  However, after return to trailer, getting bad data
from TRH:  both T and RH reading > 200 and s/n reporting as 704.

Steve cleaned KH2O at 30 m (~9:55 am) and 5 m (~10:30 am).
Thus on tower from ~9:45 am until 10:40 am PST.

Apparently, cleaning the KH2O's decreased the voltage at 30 m from 1.5 V to 
1.3 V!

54: HotFilm, Site central, Wed 15-Mar-2006 08:49:48 PST, hotfilm analogs died
cockpit shows that the A/D process on the hotfilm system died sometime between
John and I leaving at about 7 last night and us arriving about 8:45 this 
morning.  I'll leave it going, since the sonics (only the lowest 2 are enabled)
are still running.

53: power, Site central, Tue 14-Mar-2006 18:46:37 PST, cylinder readings
When we connected the fuel cell to the full cylinder yesterday morning, it was
reading 2500psi.  This morning, the reading was 2260psi -- a normal change,
for once!

52: TRH, Site south, Tue 14-Mar-2006 15:54:05 PST, 15 m TRH failed at south site
15m TRH failed at south site this morning at 8:20 am local time, March 14.
It had also been bad (210 deg = overrange?) for some previous period, at least
since 00Z 3/12.

Brad and Tom visited site ~3 pm and recycled power to no avail.

51: soil, Site west, Tue 14-Mar-2006 12:05:22 PST, TP01 wiring fixed
The wire from C7 to +12SW control never was installed in this logger.  Have
done that (and reseated the TP01 cable wires that we removed last night).
Will look at the data after a few hours and see if this works.

Connected the TPO1 that was originally installed, though the spare is still
in the ground.

Also removed the beacon control circuit for debugging.

Also gave a tour to one of the lidar ladies and took photos of the KingAir.

P.S. Didn't work.  Visited again about 18:00 (to replace the beacon circuit,
which now got it running) and found that the regulator ground pin had broken
off.  Have just fixed in the trailer (18:45) and will put back tomorrow.

50: radiation, Site central, Tue 14-Mar-2006 11:57:26 PST, Tcase working at central
After several conversations, Steve S/John/I found a programming error in the
central 4-comp logger.  The P7 instruction needed single-ended channel 3, not
2.  Made the change manually (*7h*1...*0) and cycled power to check that it
was saved.  Now Tcase ratios are 1.35, 1.34, and 1.34 for a1, a2, and 4-comp
"out"s, respectively!  The 10m values are different, but perhaps this is 
real and/or a calibration issue.

49: , Site south, Tue 14-Mar-2006 09:25:28 PST, South service visit
Brad and Tom visited south site March 14, after lunch.
Replaced two of three batteries;  John and Brad returned in the evening to
replace the third (waited until had field console to shut down NIDS).
48: , Site west, Tue 14-Mar-2006 09:14:11 PST, West service visit
Brad and Tom visited west site before noon on Monday Mar 13.  Brad and John 
returned in the evening to finish some tasks.

Brad and Tom took two new batteries to site; replaced auxillary battery;
later in day Brad and John replaced main battery (waited to change until 
had console to shut down system first).

Fixed beacon daylight sensor by squeezing terminal on red wire to fit more
tightly on pin.  Measured proper output voltage with sensor covered.
However, John reported that it was not working when they returned to the
site that evening.

Brad tripped on loose coil of wire between NIDS and logger wireless receiver.
Beeped out cable ok but connector broken on receiver end.  John and Brad 
replaced cable later in the evening(?)

Removed internal logger TP01 cable for check out at base.  Steve O. checked
regulator okay.  Brad and John replaced cable in the evening.
White H7
Green L8
'thick' Black AG
Red H9
'thin' Black L10
power: red SW12V
       black Gnd

47: radiation, Site central, Mon 13-Mar-2006 12:21:25 PST, wiring error
Found that the wire logger L2 to Amp pin 6 had been inserted in pin 8 by 
mistake.  Fixed this about 11:30.  Didn't seem to help the Tcase data :(

46: LOG, Site all, Mon 13-Mar-2006 09:13:01 PST, Daily status
ISFF Daily status 12 Mar 2006 [written next morning]

Weather: Clear and cold in the valley.  Occasional snow flurries in the \
Sierra foothills.  IOP#4 to start late Monday.

                Central         South           West     HotFilm    Base
Networking: [all up]
                ok              ok              ok       ok         ok
Data systems: [all up and archiving data on site]
                ok              ok              ok	 ok [but analog dies]

Sonic anemometer [need boom angles shot]
30m             ok              ok              ok
25m             ok              ok              ok
20m             ok              ok              ok
15m             ok              ok              ok
10m             ok              ok              ok
5m              ok              ok              ok

Krypton hygrometer [last cleaned]:
30m             5 Mar           5 Mar           4 Mar
5m              5 Mar           5 Mar           4 Mar

30m             ok              ok              ok
15m             ok              ok              ok
5m              ok              ok              ok

Rad [Rnet,, Rsw.out,, Rlw.out] [need to clean south]
                Tcase calib off ok              ok
Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
                ok              ok              TP01 bad [logger cabling?]

[Sensors unique to specific sites:]


Licor [h2o, co2]

Aerosol [not intalled, awaiting return from manufacturer]

Hot-film sonic anemometers [all deployed Sunday]
3.5m            ok
2.5m            ok
1.5m            ok

Hot-films [mounted, but working on motor control]
3.5m            X
2.5m            ok!
1.5m            X

Power [all stations stayed up (but central died Sat/Sun night.]
source:         solar+H2           solar           solar


A big hotfilm day.  All 3 hotfilm systems and sonic anemometers installed.
Middle probe started and collected data.  Top probe installed and briefly
collected data until motor crunched it -- now there are no spares.  Data system
is losing ~10% data over the network.  Hopefully data being archived at the
site, though no way to tell at the moment.  Analog data collection (the 
hotfilms) also dies sometimes -- have had to manually monitor and reset
data acquisition.

In other news:
- Added replaced 5 old batteries by 7 new batteries at central, which kept it up
over Sun/Mon night.  There is no evidence that the fuel cell is being used
(isn't using H2, isn't releasing water, central died even with fuel cell 
"connected" Sat/Sun night).  Reason unknown.
- Have added beacon photocell switches to all 3 sites to save power, but west's
didn't work.  We left it still continously powered on (aux2).
- Still don't know why Tcase is strange at central.  Voltages reported by the
logger are signficantly different, so it would appear to be a cabling problem.
This is the most serious [standard] data problem, since it causes us to not 
have 4-component radiation (for the energy balance) or longwave radiative
flux divergence data.  We still have a (lower-quality) net radiometer at 
central for the energy balance. 4-component systems at south and west are ok.
- West TP01 still bad.  Must be bulkhead cabling inside logger box.  Will try
to fix today [mon]

45: data-system, Site central, Sun 12-Mar-2006 16:51:17 PST, central died last night
Central died last night presumably due to power (we added the load of a 
second etherant plus the prometheus yesterday evening).  Came up by itself in
the morning -- probably about 1 hour lost.  Also rebooted during the day 
when new batteries (which hopefully will help the power problem) were installed
at about 1pm(?).  It appears that the local usb kept up with these cycles.

As usual, lost 15m sonic data until manually recycled both times.

44: LOG, Site south, Sun 12-Mar-2006 16:44:19 PST, south visit
visited south ~3:30-4:15:

guy wire tensions measured.

TRIME: 4.0/2.0/2.5%

Radiometers cleaned first thing.

Installed beacon switch onto aux1.

43: HotFilm, Site central, Sun 12-Mar-2006 16:37:43 PST, hotfilm stuff installed
Spent much of today at central (John/Greg/ sometimes Steve&Brad) installing
the hotfilm stuff.  Left it with middle (nominally 2m) wire installed, but
motors not working at about 3:10.

Also checked tensions -- somewhat low.  We need to retension sometime.

42: LOG, Site west, Sun 12-Mar-2006 16:34:05 PST, west visit
Visited west ~9:30-10:10.  

Tried to fix TP01 -- don't yet know if I did anything.  Only real change was
reseating +12 SW control wire.

Installed beacon controller, but didn't work.  Kept plugged into (unswitched)

Cleaned radiometers, though somewhat cloudy.

41: LOG, Site all, Sat 11-Mar-2006 19:46:38 PST, Daily status
ISFF Daily status 11 Mar 2006

Weather: Clear in the valley/snow flurries in the Sierra.  No IOP until at
least Monday night.

                Central         South           West     HotFilm    Base
Networking: [hot film network now running]
                ok              ok              ok       ok         ok
Data systems: [all normal and archiving data on site]
                ok              ok              ok

Sonic anemometer [need boom angles shot]
30m             ok              ok              ok
25m             ok              ok              ok
20m             ok              ok              ok
15m             ok              ok              ok
10m             ok              ok              ok
5m              ok              ok              ok

Krypton hygrometer [last cleaned]:
30m             5 Mar           5 Mar           4 Mar
5m              5 Mar           5 Mar           4 Mar

30m             ok              ok              ok
15m             ok              ok              ok
5m              ok              ok              ok

Rad [Rnet,, Rsw.out,, Rlw.out] [all need to be cleaned]
                Tcase calib off ok              ok
		cleaned 3/11
Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
                ok              ok              TP01 replaced

[Sensors unique to specific sites:]


Licor [h2o, co2]

Aerosol [not intalled, awaiting return from manufacturer]

Hot-film sonic anemometers [will be installed tomorrow!]
3.5m            X
2.5m            X
1.5m            X

Hot-films [will be installed tomorrow!]
3.5m            X
2.5m            X
1.5m            X

Power [all stations stayed up]
source:         solar+H2           solar           solar


Things mostly looking up!

Serviced central -- cleaned rads, checked power, checked logger wiring.  All
seem ok.  However radiometer Tcase values are still strange.  It isn't (just) a 
calibration issue, since the voltages are also quite different 
(e.g. 1.1/1.6/1.7 V).  This is our biggest data puzzle at the moment.

Installed new TP01 at west.  Didn't seem to help -- must be a logger wiring

Attended my first operations meeting!

Back to central late in day after John arrived.  Installed hot-film data system
and network (also associated laptop in base).  Now just need to hang the 
anemometers themselves.

John also installed the beacon photodiode switch at central, which should 
save us a bit of power.  Will do the other sites tomorrow.

P.S. Thanks for the rest of the care packages via John, folks!  (But why a
painted red and stencilled box??)

40: soil, Site west, Sat 11-Mar-2006 11:17:03 PST, visit
popped in at west ~10:40 to install new TP01.  This should tell us if the
problems are with the sensor or logger.  If it works, I think I did a good
enough installation to use (somewhat moist soil), though it will take a bit
to settle in.

39: radiation, Site central, Sat 11-Mar-2006 11:12:54 PST, visit
Was at central ~10:15 to:
- check fuel cell: 13.45V/1260psi
- clean all rads (including 10m for the first time)
- TRIME: 38%/11%/28% (from west to east) -- basically unusable.
- check logger cabling.  Only difference between a2 and 4-comp on K&Z is white
wire goes to "AG" on a2 rather than "G" on 4-comp.  TP01 wiring looks the
same as I remember at west.

38: power, Site central, Sat 11-Mar-2006 09:07:24 PST, added fuel cell
Added the 60W to central (with rather jury-rigged cabling) at about 18:00 last
night.  Didn't seem to do much, i.e. didn't seem like the controller was 
putting a load on it, even though it was dusk.  Cylinder pressure ~1250.  Fuel 
cell voltage output 15.55V.

We'll go out now and check up on it.

37: LOG, Site all, Fri 10-Mar-2006 19:16:36 PST, Daily status
ISFF Daily status 10 Mar 2006

Weather: Still windy most of the day.  IOP#3 ended.

                Central         South           West            Base
Networking: [our problems appear now to be solved]
                ok              ok              ok              ok
Data systems: [all normal and archiving data on site]
                ok              ok              ok

Sonic anemometer [need boom angles shot]
30m             ok              ok              ok
25m             ok              ok              ok
20m             ok              ok              ok
15m             ok              ok              ok
10m             ok              ok              ok
5m              ok              ok              ok

Krypton hygrometer [last cleaned]:
30m             5 Mar           5 Mar           4 Mar
5m              5 Mar           5 Mar           4 Mar

30m             ok              ok              ok
15m             ok              ok              ok
5m              ok              ok              ok

Rad [Rnet,, Rsw.out,, Rlw.out] [all need to be cleaned]
                Tcase calib off ok            ok
Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
                ok              ok              TP01 bad

[Sensors unique to specific sites:]


Licor [h2o, co2]

Aerosol [not intalled, awaiting return from manufacturer]

Hot-film sonic anemometers [not installed, waiting for hotfilms.]
3.5m            X
2.5m            X
1.5m            X

Hot-films [not installed, John drives them out Friday; will install Sun]
3.5m            X
2.5m            X
1.5m            X

Power [all stations stayed up]
source:         solar+H2           solar           solar


Found that the two single radiometer loggers had died overnight (one probably
died earlier) due to drained batteries.  Swapped batteries and revived.  
Unknown why the lowest drain systems failed before the other 2.  Also 
unknown why they drained in the first place with solar panels.  Old data show
that a2 died several days ago and a1's battery has been getting weak.  Perhaps
a1&a2's batteries just started out with poor charge.

Tested 60W fuel in trailer -- seemed fine, except for one pair of plates.  Have
reinstalled at central as a second input.  Rather strange cabling should be

Checked TPO1 wiring at west.  Seemed fine, no sign of cable damage. Data show
that this has never worked.  A new probe is on the way with John.  We will
also compare the logger wiring to central and/or south's.

36: radiation, Site central, Fri 10-Mar-2006 11:01:39 PST, Radiometer Power Issues
Steve and I spent this morning trying to figure out why two of the radiation stands at central and the rad stand at west were not working.  A simple fix to get them running while we troubleshoot the problem was to switch out the batteries with ones we had in the trailer.  There was a decent amount of dust on the solar panels at central, but we are unsure how much of a factor it is.  
35: base, Site base, Fri 10-Mar-2006 08:35:26 PST, dsm crash
The dsm process crashed again last night at 1720 local (actually while we were 
here, but we didn't catch it :( ).  This probably was associated with a 
TPOP crash, of which we had lots yesterday.  All stations stored data locally.

34: LOG, Site all, Thu 09-Mar-2006 17:58:40 PST, Daily status
ISFF Daily status 09 Mar 2006 [

Weather: Quite windy most of the day (from the west).  IOP#3 going on.

                Central         South           West            Base
                ok              ok              ok              ok
                                                                (now have internet!)

Data systems:
                ok              ok              ok

Sonic anemometer [need boom angles shot]
30m             ok              ok              ok
25m             ok              ok              ok
20m             ok              ok              ok
15m             ok              ok              ok
10m             ok              ok              ok
5m              ok              ok              ok

Krypton hygrometer [last cleaned]:
30m             5 Mar           5 Mar           4 Mar
5m              5 Mar           5 Mar           4 Mar

30m             ok              ok              ok
15m             ok              ok              ok
5m              ok              ok              ok

Rad [Rnet,, Rsw.out,, Rlw.out]
                10m look odd    ok              Rlw looks odd
		maybe dirt?			maybe real?
Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
                ok              ok              ok? [check TP01 again]

[Sensors unique to specific sites:]


Licor [h2o, co2]

Aerosol [not intalled, awaiting return from manufacturer]

Hot-film sonic anemometers [not installed, waiting for hotfilms.]
3.5m            X
2.5m            X
1.5m            X

Hot-films [not installed, John drives them out Friday; will install Sun]
3.5m            X
2.5m            X
1.5m            X

source:         solar           solar           solar
                (all stations stayed up)


Our final day playing with network issues.  Moved etherant at the school to get
a somewhat better signal (8, rather than 4db).  Replaced router with earlier
Dlink and got internet service to entire .14 subnet.  Installed an etherant
at the base to get the .14 subnet on an 8-port hub here.  Changed aster to
connect to this .14 subnet hub.  Gave an etherant configured for .14 to ASU.
Their internet connection now works.  Reconfigured the REAL etherant for .14
as well.  Thus, everyone should be able to connect to the internet from 
Independence.  We haven't yet tested connections originating from outside.

Still have TPOP crashes (about 6 today).  Despite these, and other network
outages while we were changing things today, all stations stayed up, recording
data locally on site.
33: LOG, Site all, Thu 09-Mar-2006 17:57:45 PST, Daily status
ISFF Daily status 08 Mar 2006 [sent by e-mail yesterday, but only just now
entered into logbook.]

Weather: Mostly clear, strong winds in the morning changed to light
in the afternoon.  IOP#3 begins tomorrow

                Central         South           West            Base
                ok              ok              ok              ok
                                                                (but no internet)
                [see COMMENTS below]

Data systems:
                ok              ok              ok

Sonic anemometer [need boom angles shot]
30m             ok              ok              ok
25m             ok              ok              ok
20m             ok              ok              ok
15m             ok              ok              ok
10m             ok              ok              ok
5m              ok              ok              ok

Krypton hygrometer [last cleaned]:
30m             5 Mar           5 Mar           4 Mar
5m              5 Mar           5 Mar           4 Mar

30m             ok              ok              ok
15m             ok              ok              ok
5m              ok              ok              ok

Rad [Rnet,, Rsw.out,, Rlw.out]
                ok              ok              ok

Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
                ok              ok              ok? [check TP01 again]

[Sensors unique to specific sites:]


Licor [h2o, co2]

Aerosol [not intalled, awaiting return from manufacturer]

Hot-film sonic anemometers [not installed, waiting for hotfilms.  Broken
        sonic appears to have revived in the base, also have spare from
        Inyo county]
3.5m            X
2.5m            X
1.5m            X

Hot-films [not installed, John drives them out Friday; will install Sat/Sun]
3.5m            X
2.5m            X
1.5m            X

source:         solar           solar           solar
                (central stayed up overnight, new batteries coming with John)


A MUCH better day.  Central did not lose power overnight.  Worked on networking
mostly.  Problem with AP24 was a change to the wrong WiFi encryption type
(should have been WEP) -- somehow we missed this yesterday.  Brought the AP24
down from the tower to work on it, made the fix, and reinstalled (now with
all CAT5 cable, so its wired interface should work at 100BaseT).  Had to visit
south to power cycle its etherant.  With this, had WiFi LAN working again at
14:30.  All stations appear to have saved their data locally during this network
outage, so no data were lost.

Continued to work on the school connection.  Managed by the end of the day to
have systems on the .14 subnet able to name-serve over the internet, but
couldn't get much else working.  Presumably, the WiFi connection is still weak
enough (4db) that significant traffic can't make it through.  We will try a
last time to adjust the antenna tomorrow, but our Boulder teammates have begun
exploring options (DSL/cable to the base).  The local WiFi contact has offered
to get REAL connected for $1700 at high-bandwidth.  It is unknown if we would/
could add on to that connection.  We also have had a routing problem that prevents
systems on the .12 subnet (e.g. the aster base computer) from being able to
use the school connection, but I think we have a workaround to connect "aster"
directly to the .14 subnet.

Also (with John's help) was able to begin testing the "bad" sonic that
was giving 8000 values when at 30m.central.  It started working fine
in the trailer, and "??" reports that it thinks it is the correct
serial number (0672).  Perhaps it was just a dirty connection that one
more reseating fixed?  I'll go through John's cheatsheet tomorrow to
verify the version4 code download and do an even better cleaning of
the connection.  In the meantime, Greg has borrowed another CSAT3 from
Inyo County that could be used as a spare.  Thus, we now have all the sonics
we need for the hot-film system when it arrives.

A note for Gordon: we have had to restart the dsm process on "aster"
(/etc/init.d/nids restart) about once a day for the last 3 days.  Presumably,
the flaky network has caused it to crash?  I also thought I found a correlation
with TPOP crashes, but at least one exception to my theory has occurred.

32: wireless-network, Site base, Thu 09-Mar-2006 10:23:26 PST, another TPOP crash
Another TPOP crash, requiring a manual reboot, just occurred.  This time 
everybody came back up (though west took a while).

31: wireless-network, Site base, Thu 09-Mar-2006 08:49:43 PST, TPOP crash
The TPOP died at 05:15 UTC (9:15PM last night), so there are no local data from
any stations until I rebooted the TPOP just now.

central is down, though the etherant is okay now.  Will have to visit it.

Both south and west recorded data locally through this outage.

P.S. central was working except for the network.  "reboot" brought everything
(including the network) back up.  All data were recorded on the usb, so none
were lost during this network outage. (Also, power stayed up overnight.)

30: LOG, Site all, Tue 07-Mar-2006 18:05:41 PST, Daily status
ISFF Daily status 07 Mar 2006

Weather: Partly cloudy, cold&windy late in the day.  IOP#2 ended 03Z (last 

		Central		South		West		Base
		ok		down		down		up
								(but no internet)

Data systems:
		up		on		on

Sonic anemometer [need boom angles shot]
30m		ok		ok		ok
25m		ok		ok		ok
20m		ok		ok		ok
15m		ok		ok		ok
10m		ok		ok		ok
5m		ok		ok		ok

Krypton hygrometer [last cleaned]:
30m		5 Mar		5 Mar		4 Mar
5m		5 Mar		5 Mar		4 Mar

30m		ok		ok		ok
15m		ok		ok		ok
5m		ok		ok		ok

Rad [Rnet,, Rsw.out,, Rlw.out]
		ok		ok		ok

Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
		ok		ok		ok? [check TP01 again]

[Sensors unique to specific sites:]


Licor [h2o, co2]

Aerosol [not intalled, FedExed today to manufacturer to set to 1 samp/s]

Hot-film sonic anemometers [not installed, awaiting hot-films/one is broken]
3.5m		X
2.5m		X
1.5m		X

Hot-films [not installed, John drives them out tomorrow; will install Sat?]
3.5m		X
2.5m		X
1.5m		X

source:		solar 		solar		solar
		(has died in early AM after clouds -- just central; new 
		batteries coming with John)


Central died last night (probably low battery).  15m sonic didn't come back up.
usb drive didn't come back up (but we were archiving over the WiFi network).
Visited station and reset drive.  Reset sonic using rserial.

Rest of the day spent trying to get school network up.  Put 180deg antenna at
west on AP24 external port.  Was able to get south, but only 12db, rather than
40db using parabolic antenna.  (Also not as much as the 20db I got yesterday 
with a roaming etherant.)  Eventually removed one RG cable which may have 
increased S/N by ~2db.

At school, redirected etherant to point to west, configured to AP24, .14 subnet.
Only 4db (eventually 5-6db).  Swapped etherants, changed RF channels, no change.

During RF channel change tests, AP24 int(!) went down, all settings seem 
reasonable as viewed with either WWW interface or winbox, but have not been 
able to revive the connection back to TPOP at base.  Thus, at the end of the 
day, can't ping west or south.

Overall, a change for the worse on all fronts!

P.S. While at west and central, checked guy wire tension (all somewhat lower).
Also took TRIME measurements at west (3.7-5.0%).  Also, played with whether
the sonics can sense someone on the tower by deliberate twisting of the tower
30 times at 11:27AM.  The RF link probably was down then, but hopefully the
usb disk has a record of this.
29: wireless-network, Site base, Mon 06-Mar-2006 12:20:31 PST, school connection weak
The school connection has been very weak.  I spent an hour on their roof this
morning (in part to help DRI with their 915 Freewave problems).  The best S/N
I could get was 4-6db -- pretty lousy.  Sometimes I can now ping, sometimes not.
Average is ~30% packets get through.

I'm convinced that I need to reconfigure the network to put the TPOP at west.
Guess we'll work on that tomorrow (when Brad is back).

Prior to this, I implemented Gordon's suggestion to lower the RTS/CTS packet 
size from 500 to 32 bytes, but of course this didn't fix school's problem.

28: power, Site central, Mon 06-Mar-2006 12:17:52 PST, central power
Central stayed up overnight.  Today is overcast and some drizzle.
Measured all 5 batteries (disconnected from each other, though 1 was kept tied
to the station)
1	12.32
2	12.41
3	12.43
4	12.42
5	12.45 (the one I added yesterday)
All seem reasonable, though Kurt thinks they should be stronger and is planning
to get new batteries to us sometime.

27: sonic, Site south, Sun 05-Mar-2006 16:57:44 PST, south 30m now serialized
With the serializer swap to CSAT 1117 mentioned earlier, we had a sonic that
could go to 30m.s.  We have just installed it.  That means, for the first time
for TREX, all standard sensors finally are working!  (Only 5 days and 1 1/4
IOPs late.)

Naturally, the network has started being flakey just to spite us.  My guess
is that the wind is moving the antennae around (though the TPOP looks quite
stable at the base).  The usb drives seem to be working through this.  I've
just played around with TPOP channels.  Gordon had left it as channel 11, but
I'm now running on 8 (after trying 2,3, and 5).

26: wireless-network, Site base, Sun 05-Mar-2006 16:29:40 PST, strange stuff
lots of network outages now/dropped data.  have changed TPOP to chan 2 and 
rebooted several times. ?????

25: sonic, Site south, Sun 05-Mar-2006 13:28:56 PST, Serializer WARNING!
Since our now 2 spare sonics consisted of a dead sonic with a serializer and
a good sonic without a serializer (and we need a serializer still at 30m.s),
We just pulled the serializer out of 672 and installed it in 1117.  Note that
this breaks the sequence that sonics with S/N <700 had serializers and newer
ones didn't.  Someone will have to make a decision later as to whether to 
swap this serializer back.

24: sonic, Site all, Sun 05-Mar-2006 12:48:12 PST, sonic S/N
After all the swapping, here is the latest list.  [Note that 20m.c was 671, not
674 as in Chris's notes.]

		Central		South		West
30		671		1119		540
25		1123		674?		1120
20		855		1122		0800
15		856		1124		539
10		832		1121		536
5		538		537		677

"Spare" [waiting to be used with the hot-film system]
672 [has serializer, but reports 8000 on all data channels
1117 [works, but no serializer]

Now all but South 30m has a serializer, as originally intended.  We will try
to put 672's serializer in 1117.

23: LOG, Site all, Sun 05-Mar-2006 12:40:57 PST, site visits
Spent the morning at all 3 sites:

Central ~9:00-10:15:
- Tried and failed to get good data from 30m sonic
- Replace 30m sonic with what was removed from 20m yesterday
- Cleaned krypton at 5m at 10:02 (since read <0.2V in the data files last 
night).  Note that all other kryptons had been cleaned when swapping sonics
with serializers.

South ~10:30-11:15:
- Swapped out 15m sonic
- Swapped in 5m sonic
- checked that all solar panels were connected

Base ~11:30:
- Restarted sonics at 5m and 15m at central via rserial [shouldn't have needed
- Restarted sonic at 15m south via rserial [also shouldn't have needed]

West ~12:00:
- checked that all solar panels were connected

Central ~12:15: (simultaneous with West visit, using 2 vehicles)
- installed 5th battery box, since Central has been dying in the early morning
the last 2 days (at least).

22: LOG, Site all, Sun 05-Mar-2006 12:25:04 PST, Procedures
A mini-cheat sheet:

- log in [if needed]: "aster", 
- check bison "Terrabeam navigator" for all known etherants+AP24 okay
- scan /var/log/messages + /var/log/isff.log
- "data_stats sock:localhost" check for all channels, data rate, packet size
- ls -l /data/projects/TREX for latest file size increase
- if needed, edits to $ISFF/projects/TREX/ISFF/ops/ops0/ads3.xml and same 
edits in ops1/ads3.xml, since some programs still use ops0 [a bug]
- "ps -ef | grep dsm" should have a "dsm_server" process running
   if not, run "/etc/init.d/nids restart"
   if this does not work, then perhaps all wireless links are down
- "ps -ef | grep stats" should have a "statsproc sock:localhost:30000"
  running to create covars and write to 
  where here "403" is April 3.  If not, see logbook entry 138

Important files/directories:
- $ASTER=/usr/local/src/isff/aster
- $ISFF=/usr/local/isff
- $ISFF/projects/TREX/isff/ops has two directories, ops0/ and ops1/, each
	of which contain ads3.xml files with the new-style project config.
	Make any changes to both (until further notice from Gordon).
- /var/log/messages and /var/log/isff.log have error messages
- /var/named has ethernet address in files 192.168.12, 192.168.14,
	that may need to be modified.

[Toughbook booting:]
- Hold power lever down ~2 sec
- After ~20 sec, down-arrow 4 times to "win2000"
- After a while, upon prompt ctl-alt-del to get login window
- log in: "lroot", 
- After another while, when win settled, click on "ProcommPlus"
- After yet another while, click on "19200" on bottom title, select "115200"
- hit  in main window 
[Now talking to viper:]
- log in: "root",
- "ps -ef | grep dsm" [should be 8 processes]
- "ping" [the ISFF base, via WiFi]
- "df" [should see 2 disks, including /var/tmp/usb]
- "rserial /dev/ttySxx localhost" [to check sensors], but often crashes nids, 
so follow by "/etc/init.d/nids restart"
- "exit" to logout upon leaving station
[Toughbook shutdown:]

GPS: (what I've been doing)
- insert memory stick
- click on "TREX" icon
- copy day directories, e.g. "062" to memory stick
- take memory stick to hotel, start VPN, scp memory stick files to 
- when memory stick full, follow instruction sheet to copy to CDR, then erase.

New GPS transfer method, (replacing last 2 steps above):
- plug memory stick in lower USB port on aster
- when window pops up with message about LEXAR_MEDIA, click OK
The copy files for day 'xxx', compare sizes (esp. if it's already on the
data disk and asks you whether to overwrite that.  Note that this is
because sometimes a day appears missing on the gps, and the gps machine
probably put it in the previous, or next, day's directory),
finally remove the files from usb stick, and umount it.
- cp -r /media/LEXAR_MEDIA/xxx /data/projects/diff_gps
- ls /media/LEXAR_MEDIA/xxx /data/projects/diff_gps/xxx
- rm -r /media/LEXAR_MEDIA
- umount /media/LEXAR_MEDIA
At midnight, the files are copied to /scr/tmp/oncley/TREX/indep_gps.


- look at terrabeam network status [laptop on shelf]
- ping central, ping west, ping south
- data_stats sock:localhost [on aster]
- data_dump -d 1 -s xx -p sock:localhost [on aster], using sid+1(+2+3+4) shown
	by data_stats
- look at cockpit, check_aster [on ripple for hotfilm]

- backup copies of files: [sometime after 00Z of day xxyy]
	cd /data/projects/TREX
	cp isff_2006xxyy* /mnt/exthd8/projects/TREX
	cp isff_2006xxyy* /mnt/exthd9/projects/TREX
	df . [if >90%, delete a few days with: rm isff_2006aabb* -- obviously,
		double check that these files are on exthd8&9!!]

- copy uncopied GPS files to memory stick, network copy to Boulder? 

Station visit [weekly]:
- clean kryptons
- clean radiometers
- TRIME soil (and gravimetric?)
- check guy wire tension and adjust if needed 
    [log in $ASTER/projects/TREX/doc/TowerTensions.xls]
- check batteries in a1, a2 loggers at central, replace if needed (and recharge
	old ones in the base) [may not be needed if Steve changes logger
- check H2 level at central, change cylinders if needed.

21: sonic, Site all, Sat 04-Mar-2006 17:23:35 PST, Sonic Update
Today consisted of relocating sonics/sonic boxes to put them in the correct positions.  The west site is completed with sonics containing serializers at 5m and 30m.  Central is completed with the exception of the 30m sonic.  We are still troubleshooting as to the data we are receiving from 30m.  We were unable to complete central due to winds.  South needs both serialized sonics mounted possibly tomorrow, pending weather conditions.  

The central site crashed again overnight last night so the morning was spent debugging the issue.  The base computer was not recognizing the DHCP and we were having issues with the USB drive.  We are still uncertain as to the problem of the central site going offline at night but the USB seems to be functioning properly at this time.

The logger cable for the NW radiometer at the central site was also repaired this morning.
20: LOG, Site all, Wed 01-Mar-2006 19:25:14 PST, Daily status
ISFF Daily status 01 Mar 2006

Weather: Clear today.  IOP started this afternoon.

		Central		South		West		Base
		ok (E'ant x1)	on (E'ant)	ok (AP24)	ok (TPOP x1)

Data systems:
		on		on		on

Sonic anemometer [need boom angles shot]
30m		reports 0	on/no kh2o	on/no kh2o
25m		on		on		on
20m		on		on		on
15m		on		on		on
10m		on		on		on
5m		on/no kh2o	on/no kh2o	on/no kh2o

Krypton hygrometer [last cleaned]: (but none connected to working serializer!)
30m		1 Mar		1 Mar		1 Mar
5m		1 Mar		1 Mar		1 Mar

30m		ok		ok		ok
15m		ok		ok		ok
5m		ok		ok		ok

Rad [Rnet,, Rsw.out,, Rlw.out]
		dirt not up	ok		ok

Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
		ok		ok		ok

[Sensors unique to specific sites:]

		ok [did 10 tip test today]

Licor [h2o, co2]

Aerosol [not intalled, needs manufacturer help to increase reporting rate]

Hot-film sonic anemometers [not installed, awaiting hot-films and data system]
3.5m		X
2.5m		X
1.5m		X

Hot-films [not installed, being built/tested in Boulder]
3.5m		X
2.5m		X
1.5m		X

source:		solar 		solar		solar
voltage:	13.24V 1 Mar	13.5V 26 Feb	13.37V 1 Mar


Started as a maintenance day, with cleaning of all rads and kryptons and
TRIME measurements (also most station voltages).  Isolated no data from 5m
sonic at central to shorted cable and blown fuse.  Replaced both.  (Now have
cable to be fixed.)

After lunch, we started looking at data.  30m sonic at central is reporting
0s (but does have a serializer).  15m sonic at central was in a strange state,
but was recoverable.  All other sonics were ok.  

Also found that one of the central radiation boxes (the bare dirt pyg) data 
are not being received.  Power seems ok.  We are guessing that it is a problem
with the Maxstream radio configuration.

HOWEVER, then looked at message lengths to check serializer positions.  Only
30m central (the bad sonic!) was correct.  2 serializers are even back in the
trailer now (they were awaiting use by the hot-films).  We are going to have
to do a massive exchange of pulling down sonics and swapping them (both heads
and boxes).  Fortunately, we still have the hot-film sonics to work with as
well.  This will require a full day to accomplish (with light winds).  Greg
and Vanda have agreed that this will be put off until the end of this IOP.

Even with all this, we may still end up with one bad sonic.  We do not have
any spares for this project, so this will have to be fixed.

Networking is now set up as we intend to be used for the rest of the project
with exceptions:
1. Sometimes our main access point dies for no apparent reason.  We changed
radio frequencies yesterday, which seemed to help, but it just died again.
2. We want a different cable at West to increase data throughput and/or
recovery from power cycles.  One has been purchased.
3. Etherants are now distributed to REAL and DLR.  REAL's has been installed.
Data systems are now operating, though perhaps not with the final code.  Data
are being archived both locally and over the network.

Station power using solar panels appears to be adequate (given that we had a
few cloudy days).  Work on getting the hydrogen fuel cells running remains
at a low priority.

19: LOG, Site west, Wed 01-Mar-2006 14:01:50 PST, west visit
West visit ~1230-1330

- cleaned rads
- cleaned kryptons
- TRIME readings: 11.2/7.7/9.7%
- aux1 voltage 13.37V

18: LOG, Site central, Wed 01-Mar-2006 14:00:30 PST, south visit
South visit ~11:30-12:40

- cleaned rads
- cleaned kryptons
- TRIME soil moisture 3.9/6.8/5.2%

17: LOG, Site central, Wed 01-Mar-2006 13:55:37 PST, central service
Visited central ~1000-1100
- 10 tips on rain gauge at 1002
- cleaned rads (except 10m)
- TRIME readings: gave "salinity too high", "no TDR signal", eventually got
readings of 20/23/40%.  (Use 40%??.)
- cleaned kryptons (5m is a hard reach)
- tried to level 5m csat better, but boom always droops (I suspect no pin in
- diagnosed csat 5m problem as cable with a short.  Will replace on the way
to lunch
- measured aux1 voltage 13.24V

16: sonic, Site central, Tue 28-Feb-2006 19:51:07 PST, central boom angles
Shot boom angles today, though the theodolite base wouldn't turn much -- had
to do a VERY coarse adjustment rotating the entire theodolite head.

Fortunately, much of the tower looked the same:

30m	93 deg 13'
25m	92 deg 35'
20m	92 deg 35'
15m	92 deg 35'
10m	92 deg 35'
5m	85 deg 21' [separate tower]

15: data-system, Site south, Tue 28-Feb-2006 19:31:30 PST, usb drive install
The usb drive had never been installed at South.  Started it about 19:00 -- our
first data from South :(.

14: data-system, Site west, Tue 28-Feb-2006 19:29:43 PST, usb drive restart
Visited west site ~1800-1840 to restart the usb drive (which needed fsck).
Also repointed the parabolic antenna to South, since south beacon was visible.
Also checked guy wire tensions (see earlier log entry).

13: LOG, Site all, Tue 28-Feb-2006 17:32:04 PST, Daily status
ISFF Daily status 28 Feb 2006

Weather: Partly cloudy and cooler after last night's rain.  Not as dusty.

		Central		South		West		Base
		up (E'ant x1)	up (E'ant)	up (AP24)	up (TPOP x1)

Data systems:
		on		on		on

Sonic anemometer [need boom angles shot]
30m		on		on		on
25m		on		on		on
20m		on		on		on
15m		on		on		on
10m		on		on		on
5m		off		on		on

Krypton hygrometer [last cleaned]:
30m		on		on		on
5m		on		on		26 Feb

30m		on		on		on
15m		on		on		on	
5m		on		on		on

Rad [Rnet,, Rsw.out,, Rlw.out]
		on		on		on

Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
		on		on		on

[Sensors unique to specific sites:]

		on [should see something today]

Licor [h2o, co2]

Aerosol [not intalled, needs manufacturer help to increase reporting rate]

Hot-film sonic anemometers [not installed, awaiting hot-films and data system]
3.5m		X
2.5m		X
1.5m		X

Hot-films [not installed, being built/tested in Boulder]
3.5m		X
2.5m		X
1.5m		X

source:		solar 		solar		solar
voltage:	13.1V 28 Feb	13.5V 26 Feb	13.4V 26 Feb


Boom angles shot today at Central, though theodolite not working well.  Might
have to be redone.

All standard sensors are up and all but one sonic have now been confirmed as
reporting.  We still do not have access to the data to evaluate quality.  

Networking is now set up as we intend to be used for the rest of the project
with exceptions:
1. We want a different cable at West to increase data throughput and/or
recovery from power cycles.
2. Etherants need to be distributed to REAL and DLR
Data systems are now operating, though perhaps not with the final code.  Data
should be being archived locally now.  We found that South's local storage had
never been installed, so its data set started an hour ago.  (3 hours into the
official operations period.)

Most sensors need standard service (cleaning and boom angles shot), which we've 
deliberately delayed until close to the start of operations.  We'll work on
these tasks tomorrow, in preparation for the IOP scheduled to start tomorrow 

12: LOG, Site all, Tue 28-Feb-2006 16:59:59 PST, guy wire tensions
Guy Wire tensions

CENTRAL:	Initial	28 Feb
Wind?			W@8?
4		425	330
3		450	390
2		400	350
1		560	480
4		450	330
3		400	340
2		440	440
1		525	490
4		400	290
3		400	310
2		400	350
1		520	440


SOUTH:		Initial	27 Feb
Wind			S@8?
4		340     400
3		350	330
2		310	375
1		400	400
4		310	400
3		440	400
2		300	410
1		475	480
4		350	410
3		490	450
2		340	460
1		480	480

WEST		xx Feb	28 Feb
Wind			calm
4		400	340
3		410	350
2		400	350
1		500	430
4		400	325
3		400	350
2		420	350
1		475	410
4		460	380
3		500	450
2		460	390
1		550	500

11: wireless-network, Site all, Tue 28-Feb-2006 14:40:06 PST, network stuff
Two network improvements today:

- Got school connection working (mostly by turning off IPTABLES).  Also found
that this laptop won't reboot due to some docking station power issue.  
Installed (probably very weak) battery to smooth power surges?

- Got south working as we want.  Fixed cable to top etherant, set etherant
to static IP, gateway of AP24 external, removed lower etherant.

Still want to exchange strange cabling at west to AP24 with a good cat5 cable.

10: base, Site base, Mon 27-Feb-2006 21:34:47 PST, monitoring tools
data_stats sock:localhost [gives all nids channels]

Alico AP24 winbox (on Toughbook) [gives West's status]

TeraBeam navigator/configure/click on 192...address/"configure remote"/
"AP Associated ..." (on Toughbook) [gives Etherant/AP24 internal status]

More Etherant status on their own WWW sites.  (see /etc/hosts for addresses)
[set up as pull-down bookmarks from mozilla on aster base]
passwords should be saved by mozilla.

9: LOG, Site all, Mon 27-Feb-2006 20:42:26 PST, Old status messages
[Here are the e-mail status messages sent by various people during set-up,
prior to logbook being started.]

Status, 2/7/06 [semmer]

Kurt received a call this morning from the building inspector giving
us premission to start the tower installation. So we headed to the
south tower to install the anchors. The manta rays went in without any
trouble. We completed the six anchors in about 2 hours. From south we
headed to Independence for lunch then on to the central site. At
central the first 2 anchors went in without any problem. The next two
were not as easy. We had to pull out the 88' anchor twice before we
were able to get it in deep enough. The final two went in without any
problems.  Tomorrow we hope to get the west anchors in and possibly
start installing the west tower.

Status 2/08/2006 [semmer]


Since you talked to Kurt you have the latest update. The anchors at
west gave more problems than central. we got them in but not as deep
as we wanted; however, the pull test better than 8000 lbs. If things
continue as planned I could see us having a 3o meter tower up by
tomorrow.  Steve

PS - Kurt was going to call John to about adding two more of the solar panel charger boards in the battery box. One of the boxes sent out did not have the board in it so I would like to have a few spares. 

>> Hi Steve,
>> I was very glad to get Kurt's call that the anchors were done today.  
>> That must be a relief to everyone there and I've spread the news around 
>> here.
>> The driver called and plans to show up at 8:00 AM tomorrow to take the 
>> base trailer.  John is finalizing packing of the network stuff as I 
>> write this.  I've put a little note by the drill press to point you to 
>> the location of items that might not be obvious.
>> Good luck! See you soon...Steve

Status 2/13/2006 [semmer]

Friday: Finished the installation of the south tower. The fencing was
installed and the beacon. The dark horse still needs to be located.
The base station arrived today.

Saturday: Started installation on the west tower. Got up to 30 ft.

Sunday: Continued work on west tower. Ran into some problems with the
vertical alignment. This may have been due to two sets on guy wires
installed before the we tryed to align. At the end of the day the
tower was at 80 ft.

Plans for today: Finish off the west tower including the electricl
fence and the soil sensors. If time permits start on the central

Status 2/13/2006 [semmer]

Installation of the west tower infrastructure was finished today. This
included a final transect alignment of the tower along with setting
the guy wires to the proper tension. The electric fence was
installed. Also the 60 watt fuel was setup. It worked as expected. The
hydrogen bottle was attached, gas flow started, and the fuel cell
immediately put out 14 volts. I played around with the fans to try to
improve the output. The wiring in the box was changed so that an
external battery could be charged. The beacon light was attached to
provide a load. Tomorrow I will check how it is doing. Tomorrow's goal
is the central tower. Also we will have the building inspector check
out the south and west towers.

Status 2/14/2006 [semmer]

Work started on the central tower today. We got to 90' then the wind
started blowing making it too difficult to raise the gin
pole. Tomorrow we will finish off the central main tower and start
setting up the 10 meter and 5 meter ASTER towers. Testing of the
wireless communication may start tomorrow.
The building inspector checked out the south tower today. There were
no problems.

Status 2/15/06 [semmer]

Due to high winds today we were not able to finish the installation of
the central tower. So we spent most of the day at the base trailer
getting power hooked up, organizing, and working on the wireless
communication. The communication modules were checked out a the base
then one Etherant was taken to the South tower while a PAM tripod was
setup at the base to handle the 180 degree antenna. Unfortunately we
were not able to communicate between south and base. If the weather is
good tomorrow we will try a communication test with the TPOP at west.

Status 2/17 [Golubieski]

Dressed Central 30M tower.  Found out some CSAT booms do not fit
into our new boom clamps.  Also, the Krypton electronic box mounting
bracket did not work.  Need more room to mount the box on the inside
of the tower.  We ended up mounting the Krypton in the inside of the
tower but used quicklinks and zip ties.

105feet - Bridge H
29.98M - CSAT 0672, Krypton 1394, TRH ?.
24.90M - CSAT 1123
20.00M - CSAT 0674
15.00M - CSAT 0856, TRH ?.
10.00M - CSAT 0538

Status 2/18 [Golubieski]

Dressed West tower.  Some Krypton electronic boxes mounted on the
outside of tower.  We were short one CSAT cable to finish the cabling.
Noticed the safety-line took some damage during the past few days.
Some chaffing by guy bracket areas.  Need to keep an eye on it.  We
tightened the life-line and seems to be better?.but there is no wind.
Hung Terra-Beam at ~105 feet.  Got a wireless signal but did not see
it at the base station.

105feet - Bridge G
30.00M - CSAT 1124, Krypton 1258, TRH ?
25.00M - CSAT 1120
20.00M - CSAT 0800
15.00M - CSAT 0539, TRH ?
10.00M - CSAT 0536
4.97M - CSAT 0855, Krypton 1395, TRH ?
2.00M - Barometer U4110003

Status 2/19 [Golubieski]

Poor weather today.  Snow and wind.  Found out that Central tower had
no power when arriving in the morning.  Fuel cell stopped working
(Voller).  Restarted Voller but shut down in a matter of minutes?when
I unplugged cable from battery boxes.  Voller's battery was drained
completely.  Once the battery was charged Voller shut down NUMEROUS
times showing the "NO HYDROGEN" remark.  Kept trying to keep Voller
going but was unsuccessful.  Noticed that if we unplugged all
batteries but one the Voller stayed running.  It seems that we were
putting too much load on the fuel cell.  So we brought over three
solar panels to help charge three batteries and run beacon while the
Voller charged one battery and ran NDAQ.  Ran for hours when setting
up power this way.  Tomorrow we plan on moving some batteries to the
Voller power.  Hung barometer at 2.00 meters and connected it to the
NDAQ.  Hooked up GPS and did another test with Terra-Beam antenna.
Still reading "acceptable" to "good."  Will fine-tune when weather
permits.  Began making fence around whole area of Central.  Got most
of the way with remaining fence posts but it's not complete.  Ten
posts short.  Strung barbwire on two of the five sides.  Went to West
site and hung barometer at the 2.00meter level, ran the last CSAT
(4.97M) cable to make this tower complete.  Terra-Beam test showed
"good" signal, but it could get better.  Will align when weather
permits.  It seems that our test yesterday at the base did see West's
signal but Central must have been down.  Hydrogen tank at ~800psi.

Status 2/20 [golub...]

Sunny, cold and windy.  No tower climbing in the morning.
Worked on barbwire fence around Central site and erected 10M Aster
tower.  More problems with Voller.  Did not last throughout the
night.  Set up solar panels to charge
controller?controller(w/battery) to three batteries...fused load to
NDAQ?AUX2 running beacon?end of battery chain Voller is plugged in.
In the afternoon we dressed South 30M tower.  Lots of cows to keep
them company.

***Please notice that the heights given are at the TOP of the boom clamps.***

105feet ? Bridge H
30.00M ? CSAT 1119, Krypton 1393, TRH ?
24.99M ? CSAT 0674
20.00M ? CSAT 1122
15.07M ? CSAT 0537, TRH ?
10.00M ? CSAT 1121
5.03M ? CSAT 1117, Krypton 1391, TRH ?
2.00M ? CSAT S0610003

Status 2/21:  Day off. [Golub...]

Status 2/22: [Golub...]

Finished barbwire fence at Central.  Hung K&Z and electronic box
onto 10M tower.  Did not plug in/turn on box.  Waiting to mount solar
panel.  Voller still saying, ?NO HYDROGEN.?  Will bring back to base
station and see what could be the cause.  Went back to West tower to
replace Hydrogen cylinder.  When we arrived fuel cell was not working.
Used all of hydrogen in the cylinder.  We realized that this should
not be done.  Leave a little bit of hydrogen in the cylinder.
Replaced cylinder and could not get fuel cell to charge battery.
Realized it was a big load on the fuel cell.  Battery reading 6.0V.
So had car charge battery while fuel cell ran and charged another
battery to use for NDAQ. Tried just plugging NDAQ in directly but
couldn?t keep up.  Were able to run with battery (11.8V), NDAQ and
beacon?.but not long.  Enough to get our test done for antenna
testing.  Left site with this set-up but not getting signal from
Bridge G.  Beacon is still on.  Will check tomorrow.  Steve O. and
Gordon M. arrived.

Status 2/24 [oncley]

I'm sorry to say that this is my first status report from the field.
I'll try to improve the frequency (and settle on a [more complete]
format) soon:

- All ISFF towers are up (with the exception of the short [5m] tower
at Central that will hold the hot-film and other low-level sensors).
Building inspections have been completed.

- All sensors (flux, soil, radiation, etc.) are deployed with the
exception of:
   - Two pyrgeometers needing construction of mounting plates
   - All sensors on the afore-mentioned 5m tower
   - Shane's aerosol counter

- The ISFF base trailer is operational, especially now that most
sensor boxes have been moved to the extra REAL seatainer.  (Thanks,
Shane!)  I haven't actually heard whether our official written lease
for the trailer site was finalized, but we had many assurances that
we could procede.  We have had visits by ISS (who will install a
SODAR) and Leeds (who will install their dropsonde system).  We
expect to receive the G-V GPS base station tomorrow.  All of our
hydrogen fuel cell systems have had problems and all are now in our
base for testing.  We have installed solar panels at all 3 sites,
but have not yet evaluated if these will be sufficient to keep our
stations powered throughout the project.  Work on getting the fuel
cells back on line is continuing.  We finished configuration of our
long-range WiFi network today at our base.  Now deployment of some
more parts of this network (adding a repeater to West, reconfiguring
South) need to be done.  We won't be connected to the external
network until at least Monday, due to a lack of weekend access to
the school.

We still expect to have the vast majority of our equipment operational
by the start of T-REX.  Possible or known exceptions include:
- The hot-film system, which will be constructed next week in Boulder.  
We hope it will be operational by the 2nd week of T-REX.
- WWW-pages with ISFF data may be a few days late getting online
- Networking to the LIDAR sites probably won't occur before Tuesday.

- Station power may not be quite adequate for a while.  We may delay
deployment of our Li7500 fast-response H2O/CO2 sensor until we have a
better idea of our power system capacity.

- We still need to address some basic data quality issues: sensor
cleaning, sensor alignment documentation, sensor heights, etc.  Some
of these won't get done until a few days into operations as well.

This should give you a general idea -- we are running a bit late and
still have a fair amount to do, but should have most primary data by
the time operations start.

Status 2/27 morning [oncley]

All infrastructure working.

Networking between sites (almost) working.  Network connection to the 
internet to be built today (27 Feb).

All sensors deployed except:
- hot-film sensors (and associated sonic anemometers) awaiting system 
construction in Boulder.  ETA 1 week.
- CO2 flux sensor (may be deployed today)
- Aerosol sensor in support of REAL (ETA ~10 days)

Extensive software development is ongoing.  We have no way to examine
data yet.  ETA ~3 days.

8: wireless-network, Site all, Mon 27-Feb-2006 20:10:53 PST, Network configuation
This is a stab at defining the network we've settled on:

(see /etc/hosts for complete list of machine names and /etc/dhcpd.conf for
MAC addresses)

.20 subnet (wired; in base trailer; the 8-port hub)
- aster [base computer]
- toughbook
- other users [currently Gordon's and my laptops]
aster is a DHCP server for this network

.12 subnet (wireless; access point is TPOP channel 1 at base)
- TPOP wired interface (also the 4-port hub)
- TPOP external antenna interface(?)
- Prometheus (for hot-films) through TPOP external antenna
- aster (wired to 4-port hub)
- Etherant at Central (wireless)
- nids at Central (wired from Etherant)
- AP24 internal antenna at West (wireless)
- Etherant at Independence Unified School (wired from Etherant)
- "school" laptop in the school attic [our internet gateway]
  (This laptop also on the Independence school local net, but is accessed by
  yet another address using "destination NAT" from Bishop's school.)
- Etherants at lidar sites [not yet installed] (wireless)

.13 subnet (wired interface of AP24 at West)
- AP24
- nids at West

.14 subnet (wireless; access point is AP24 external [parabolic] antenna)
- AP24 external antenna at West
- Etherant at South
- nids at South

7: wireless-network, Site south, Mon 27-Feb-2006 20:03:42 PST, South network task
The Bulgin connector on South's etherant was damaged during installation, but
we don't want to have to strip it from all the way up the tower.
At the moment, our workaround was to install another etherant at the bottom
of the tower (using rope!), since it still has good line-of-sight to West's
AP24.  This was using another, short, cable.

Brad has rebuilt the pins on the cable to the original etherant, but was 
unable to finish installing the housing and backshell due to intense blowing
dust yesterday afternoon.  On our next visit, he'll finish this task and we
can get back to a normal configuration.  The top etherant will have to be
reconfigured since it is set now to talk to the TPOP (.12 subnet), rather than
the AP24 external antenna (.14 subnet).

6: wireless-network, Site west, Mon 27-Feb-2006 19:50:44 PST, AP24 setting
We just found that we can only run at 10BaseT speeds using the cabling that we
have to the AP24 at West.  Unfortunately, the AP24 appears to go to 100BaseT
on power-up.  Thus, we have a sequence:
- power-up AP24
- run the "Alico AP24 winbox" application on the Toughbook
- connect (should be default)
- click on "ether1"
- set "10MBPS" and uncheck "Auto..." and "Full..." on "Ethernet" tab
- click on "Apply"
- ssh in from an outside machine (not necessarily easy)
- "reboot" nids

Since this is an unwieldy procedure, we need to either
- make a new cable (cat5 or cat6) that is long enough (18m)
- figure out how to save 10BaseT as an AP24 power-up default

5: soil, Site all, Mon 27-Feb-2006 19:41:26 PST, soils
Soils were installed 23 Feb at Central and South, then 24 Feb at West.  All
soils were dry and brittle the first 0-3cm, but had a moist layer below.
(Presumably from the snow that came about 18? Feb.)  Thus, I feel reasonably 
good about setting the lower sensors.  However, there definitely was an air gap
when installing especially the TP01s.  Thus, I am very glad (just from a soil 
perspective!) that we are now having a good rain.

I have not yet taken any TRIME or gravimetric samples.  I argued that it didn't
make sense to measure manually before we were recording, but this was probably
wrong -- local nids storage may have been working.  Also, I didn't have the
scales for processing gravimetric samples.  (One may be on its way from Boulder,
or I may have to find one here.)  Thus, I've missed a reasonably dry period 
for checking calibration.  Ordinarily, we wouldn't need gravimetric since the
TRIME works, but I need at least one bulk density measurement at each site.

4: LOG, Site all, Mon 27-Feb-2006 19:08:11 PST, Daily status
ISFF Daily status 27 Feb 2006

Weather: clear and mild from the time I got here (22 Feb) through day before 
yesterday (25 Feb).  Yesterday, became cloudy with occasionally strong winds
from the S and blowing dust in the valley.  Today, continued S winds with dust
and occasional drizzle.  This afternoon, drizzle increased to light rain.  This
should be great for setting the soil sensors, stabilizing some of the dust, and
perhaps cleaning some of the sensors.

		Central		South		West		Base
		up (E'ant x1)	up (E'ant)	up (AP24)	up (TPOP x1)

Data systems:
		up		up		up

Sonic anemometer [need boom angles shot]
30m		up		up		up
25m		up		up		up
20m		up		up		up
15m		up		up		up
10m		up		up		up
5m		up		up		up

Krypton hygrometer [last cleaned]:
30m		up		up		up
5m		up		up		26 Feb

30m		up		up		up
15m		up		up		up	
5m		up		up		up

Rad [Rnet,, Rsw.out,, Rlw.out]
		up		up		up

Soil [Gsoil, Tsoil, Qsoil, Cpsoil]
		up		up		up

[Sensors unique to specific sites:]

		up [should see something today]

Licor [h2o, co2]

Aerosol [not intalled, needs manufacturer help to increase reporting rate]

Hot-film sonic anemometers [not installed, awaiting hot-films and data system]
3.5m		X
2.5m		X
1.5m		X

Hot-films [not installed, being built/tested in Boulder]
3.5m		X
2.5m		X
1.5m		X

source:		solar 		solar		solar
voltage:	13.1V 27 Feb	13.5V 26 Feb	13.4V 26 Feb


Although all standard sensors are up, data are not yet available to evaluate
functionality/quality.  Most also need standard service (cleaning and boom 
angles shot), which we've deliberately delayed until close to the start of

3: wireless-network, Site base, Mon 27-Feb-2006 11:03:25 PST, school internet contacts
Independence: Joel Hampton superintendent 760-878-2405
Bishop: Joe Griego IT director 760-920-1800 (c) 760-872-7027 (office)
        Justin Norcross Network manager 760-920-8959 (c)

2: project-info, Site all, Thu 23-Feb-2006 10:24:55 MST, Phone numbers
Phone numbers: 

TREX Ops Center 760-872-4214

ISFF Eng: 	303 579-7860
ISFF Scientist: 303 319-5585

Kurt: 		303 579-1130
Brad Agee:	765 717-4844
Chris G: 	303 250-7749
Gordon: 	303 579-6026
Greg P:		303 579-7302

Renee: 		303 818 4303

Shane (REAL):  	720 938-1144
Laura (REAL): 	303 717-6382

1: LOG, Site none, Tue 21-Feb-2006 12:26:53 MST, Logbook created
initial log file /usr/local/src/isff/aster/projects/TREX/logbook/tklog.log created by USER aster